### v3

# Logit-Q Dynamics for Efficient Learning in Stochastic Teams

Abstract
We present a new family of logit-Q dynamics for efficient learning in stochastic games by combining the log-linear learning (also known as logit dynamics) for the repeated play of normal-form games with Q-learning for unknown Markov decision processes within the auxiliary stage-game framework. In this framework, we view stochastic games as agents repeatedly playing some stage game associated with the current state of the underlying game while the agents’ Q-functions determine the payoffs of these stage games. We show that the logit-Q dynamics presented reach (near) efficient equilibrium in stochastic teams with unknown dynamics and quantify the approximation error. We also show the rationality of the logit-Q dynamics against agents following pure stationary strategies and the convergence of the dynamics in stochastic games where the stage-payoffs induce potential games, yet only a single agent controls the state transitions beyond stochastic teams. The key idea is to approximate the dynamics with a fictional scenario where the Q-function estimates are stationary over epochs whose lengths grow at a sufficiently slow rate. We then couple the dynamics in the main and fictional scenarios to show that these two scenarios become more and more similar across epochs due to the vanishing step size and growing epoch lengths.
1 Introduction
Stochastic games, introduced by (Shapley, 1953), are generalizations of Markov decision processes to non-cooperative multi-agent systems and used as ideal models of multi-agent reinforcement learning in frontier artificial intelligence applications, from playing complex board games to planning for intelligent and autonomous systems (Zhang et al., 2021). Hence, the theory of learning in stochastic games has recently attracted significant interest to address whether non-equilibrium adaptation of self-interested agents would reach equilibrium, e.g., in zero-sum or identical-interest stochastic games (Baudin and Laraki, 2022; Leslie et al., 2020; Sayin et al., 2022a, 2021, b). Such results strengthen the predictive power of equilibrium analysis for games (Fudenberg and Levine, 2009). However, for control and optimization applications, a common drawback in this line of work is that they provide convergence guarantees only for some equilibrium. Agents following these learning dynamics can attain arbitrarily poor performance even in stochastic teams where every agent has a common goal. For example, efficient learning dynamics have been studied extensively for games played repeatedly (see (Bistriz and Leshem, 2021; Chasparis et al., 2013; Marden and Shamma, 2012; Marden et al., 2014; Pradelski and Young, 2012)). However, the existing results in the context of stochastic games have been very limited (Yongacoglu et al., 2022).
1.1 Contributions
In this paper, we provide an affirmative answer to the question whether we can address efficient learning in stochastic teams through non-equilibrium adaptation of self-interested agents. We present a new family of logit-Q dynamics combining the classical and independent log-linear learning dynamics (Marden and Shamma, 2012; Tatarenko, 2017) with Q-learning (Watkins and Dayan, 1992) within the stage-game framework. The family members differ in terms of when agents revise their actions and how they model their joint play, as tabulated later in Table 1. In the stage-game framework, we view stochastic games as agents are repeatedly playing some stage game associated with the current state of the underlying stochastic game while the payoffs of these stage games are determined by the agents’ immediate and discounted expected continuation payoffs, called Q-function.
The stage-game framework goes back to the introduction of stochastic games by (Shapley, 1953), also used in Minimax-Q (Littman, 1994) and Nash-Q (Hu and Wellman, 2003) algorithms to compute equilibrium in stochastic games. Recently, the stage-game framework has also been used by (Baudin and Laraki, 2022; Leslie et al., 2020; Sayin et al., 2022a, 2021, b) to combine the best response and fictitious play dynamics with value iteration for learning in stochastic games. Following this trend, in the logit-Q dynamics, we let agents follow log-linear learning in the stage games specific to each state whenever the associated state gets visited as if the stage-game payoffs are their Q-function estimates so that they can learn to coordinate in the soft maximizers of the stage games. The agents, not knowing the underlying dynamics, learn their Q-functions by evaluating the value of their play in these stage games in a model-free way, as in the (soft) Q-learning (Watkins and Dayan, 1992) as if their estimated play is the (soft) maximizer. We consider two schemes for how agents can model their past play: In the averaging approach, agents take the empirical average of the actions played. In the exploration-free approach, agents use the most frequently played actions since they play the best response with higher probability in the log-linear learning dynamics.
We show that in the logit-Q dynamics, the Q-function estimates converge to the Q-functions associated with the efficient equilibrium and the estimated play reaches efficient equilibrium in stochastic teams approximately almost surely, with quantifiable approximation errors (See Theorem 1). Furthermore, we show the rationality of the dynamics against agents following pure stationary strategies (See Corollary 1) and the convergence of the dynamics in certain stochastic games in which the stage-payoffs induce potential games while only a single agent controls the state transitions (See Corollary 2).
1.2 Challenges
A stochastic game is a multi-agent generalization of Markov decision processes (MDPs) in which agents jointly control the immediate payoff they receive and the evolution of the underlying state and, therefore, the continuation payoff. The trade-off between immediate and continuation payoffs poses a challenge for the generalization of the existing results in efficient learning for games with repeated play (such as (Chasparis et al., 2013; Marden and Shamma, 2012; Marden et al., 2014; Pradelski and Young, 2012)) to stochastic games. The immediate payoff function is stationary, whereas the continuation payoff function is not necessarily stationary by depending on the agents’ evolving strategies. Recently, the non-stationarity challenge for learning in stochastic games has been addressed via the two-timescale learning framework for the best response and fictitious play dynamics (Leslie et al., 2020; Sayin et al., 2022a, b). However, these dynamics have convergence guarantees only for some equilibrium, possibly inefficient with respect to the common team objective.
The log-linear learning update can attain the efficient equilibrium in the repeated play of identical-interest (or potential) games in the sense that the agents play the pure efficient equilibrium at high percentage of time in the long run (Marden and Shamma, 2012). Recall the challenge that the stage-game payoffs, i.e., Q-function estimates, are not necessarily stationary. Therefore, the key question is whether the action profiles played by the agents can track the efficient equilibrium associated with the evolving Q-function estimates of the agents. The two-timescale stochastic approximation methods applied in (Leslie et al., 2020; Sayin et al., 2022a, 2021, b) are not applicable to this problem as the log-linear learning update is not in the form of a stochastic approximation recursion different from the fictitious play and best-response dynamics.
We develop a new approach to address the non-stationarity challenge for the logit-Q dynamics. Particularly, we divide the horizon into epochs with growing lengths only for analysis so that we can bound how much the Q-function estimates can change and the vanishing step sizes used in the Q-function updates imply that these Q-function estimates, and therefore, the stage games become quasi-stationary within these epochs. To address quasi-stationarity, we consider a fictional scenario where the Q-function estimates do not change so that we have stationary stage games within these epochs. For the fictional scenario, the logit-Q dynamics reduce to the classical or independent log-linear learning dynamics for the repeated play of each stage game and the action profiles played by the agents form an irreducible and aperiodic Markov chain. Correspondingly, we can view the change in the Q-function estimates as perturbations for the Markov chain in the fictional scenario. However, the perturbation is not necessarily stationary or convergent, and the existing perturbation bounds for the analysis of Markov chains (e.g., see (Cho and Meyer, 2001) for an overview) are not tight enough for our analysis. Therefore, we formulate a new perturbation bound (see Lemma 1) to approximate the game dynamics with its unperturbed version by coupling the evolution of the dynamics in the perturbed and unperturbed versions. Lemma 1 can be of independent interest.
1.3 Related Works
Stochastic teams have received attention in the multi-agent learning literature. The early results mainly focused on the computation of Q-function estimates associated with efficient equilibrium in stochastic teams when the model of the underlying game is unknown, e.g., (Boutilier, 1996; Lauer and Riedmiller, 2000; Littman, 2001; Wang and Sandholm, 2002). However, these approaches are based on the maximization over action profiles in a coordinated manner rather than agents are learning to coordinate in the maximizing joint actions in an uncoupled way without taking the objective of other agents into account, as in the logit-Q dynamics for stochastic teams or as in (Chasparis et al., 2013; Marden and Shamma, 2012; Marden et al., 2014; Pradelski and Young, 2012) for the repeated play of games. Such uncoupled learning dynamics are of interest for non-cooperative environments (e.g., non-cooperative coordination problems) as they are consistent with the self-interest of the agents to a certain extent and can lead to rational response against opponents playing according to some stationary strategies.
For example, in the classical log-linear learning, $(i)$ agents believe that the others would play their latest actions (as in Cournot learning, e.g., see (Cheung and Friedman, 1997)), $(ii)$ the agents respond according to the smoothed best response to that belief when they update their actions, and $(iii)$ the agents take random turns in updating their actions. Experimental studies suggest that human behavior in games is closer to Cournot learning than the widely studied fictitious play (Cheung and Friedman, 1997). In (Marden and Shamma, 2012), the authors showed the efficiency of the classical log-linear learning and its independent and payoff-based variants in achieving the optimal equilibrium in potential games (and more generally weakly acyclic games) played repeatedly. Later, in (Marden et al., 2014) and (Pradelski and Young, 2012), the authors addressed efficient learning in general class of games beyond potential or weakly acyclic games by letting agents update (or not) depending on whether they are in the content or discontent states according to benchmark payoffs and actions they keep track of. In (Chasparis et al., 2013), the authors presented a learning dynamic where agents continue to play an action as long as the received payoff exceeds a specified aspiration level in coordination games. The common theme in the dynamics presented in these works (Chasparis et al., 2013; Marden and Shamma, 2012; Marden et al., 2014; Pradelski and Young, 2012) are recency in the beliefs formed, and exploration and friction in action revision to selectively attain the efficient outcome. Furthermore, these results are only for the repeated play of games. On the other hand, recently, fictitious play and policy gradient based dynamics have been presented for learning in stochastic teams, e.g., see (Baudin and Laraki, 2022; Sayin et al., 2022b; Leonardos et al., 2022); however, they provide convergence guarantees only for some (possibly inefficient) equilibrium.
There are a few (recent) studies relevant to the logit-Q dynamics, e.g., (Bistriz and Leshem, 2021), (Yongacoglu et al., 2022), and (Unlu and Sayin, 2023). In (Bistriz and Leshem, 2021), the authors have studied efficient learning in a special class of stochastic games in which the state transitions are not controlled by the agents. Hence, the agents do not face the trade-off between immediate and continuation payoffs, whereas this trade-off is the core challenge for learning in general classes of stochastic games. In (Yongacoglu et al., 2022), the authors focused on efficient learning in stochastic games by restricting the strategy spaces of the agents to pure stationary strategies and by dividing the horizon into phases within which the agents do not update their strategies so that they can learn the associated Q-functions based on local information only. Lastly, in (Unlu and Sayin, 2023), a precursor of this paper, we focused on episodic logit-Q dynamics, where agents do not update their Q-function estimates within the episodes, and therefore, the stage games they play are stationary (rather than quasi-stationary). However, such episodic approaches followed in (Yongacoglu et al., 2022) and (Unlu and Sayin, 2023) require the coordination of the agents in not updating certain parameters for certain time intervals periodically.
1.4 Organization
The paper is organized as follows. We describe stochastic games in Section 2 and the logit-Q dynamics in Section 3. We present the convergence result and the proofs, resp., in Sections 4 and 5. In Section 6, we provide numerical examples, corroborating the convergence results in Section 4. We conclude the paper with some remarks in Section 7.
2 Stochastic Games
Formally, an $N$-agent stochastic game, introduced by (Shapley, 1953), is a multi-stage game played over infinite horizon that can be characterized by the tuple $\langle S,(A^{i})_{i\in[N]},(r^{i})_{i\in[N]},p,\gamma\rangle$. At each stage $t=0,1,\ldots$, the game visits a state $s$ from a finite set $S$ and each agent $i$ simultaneously takes an action $a^{i}$ from a finite set $A^{i}$ to collect stage-payoffs. The formulation can be extended to state-variant action sets rather straightforwardly. Agent $i$’s stage-payoff is denoted by $r^{i}:S\times A\rightarrow[0,1]$. Similar to MDPs, stage-payoffs and transition of the game across states depend only on the current state visited and current action profile $a=\{a^{i}\}_{i\in[N]}$ played. Particularly, $p(s^{\prime}\mid s,a)$ for each $(s,a,s^{\prime})\in S\times A\times S$ denotes the probability of transition from $s$ to $s^{\prime}$ under action profile $a$, where $A:=\prod_{i}A^{i}$. Lastly, $\gamma\in[0,1)$ is the discount factor.
Assume that the agents can correlate their randomized actions and consider joint Markov stationary strategies $\pi:S\rightarrow\Delta(A)$ determining the probability of the joint actions to be played by all agents contingent on the current state. [^1] We let $\Delta(A)$ denote the probability simplex over the finite set $A$. Given $\pi$ and the initial state $s\in S$, agent $i$’s utility is given by
$U^{i}(\pi;s):=\mathrm{E}_{a_{t}\sim\pi(s_{t})}\left[\sum_{t=0}^{\infty}\gamma^{t}r^{i}(s_{t},a_{t})\mid s_{0}=s\right],$ (1)
where $(s_{t},a_{t})$ denotes the pair of state and action profile realized at stage $t$. The expectation is taken with respect to the randomness on $(s_{t},a_{t})$ induced by the underlying transition kernel and the joint strategy $\pi$.
Given a joint Markov stationary strategy $\pi$, we denote the joint strategy of agents $j\neq i$ by $\pi^{-i}$, i.e., $\pi^{-i}(s)\in\Delta(A^{-i})$ is the marginal distribution of $\pi(s)$ over $A^{-i}:=\prod_{j\neq i}A^{j}$. We also denote the joint strategy to randomize joint actions at each state $s$ according to the product distribution $\tilde{\pi}^{i}(s)\times\pi^{-i}(s)\in\Delta(A)$ by $\tilde{\pi}^{i}\times\pi^{-i}$. In the following, we describe equilibrium concepts for such joint Markov strategies.
Definition 1.
A Markov stationary strategy $\pi:S\rightarrow\Delta(A)$ is $\epsilon$-approximate perfect Markov coarse correlated equilibrium (abbreviated as $\epsilon$-CCE) provided that
$U^{i}(\pi;s)\geq U^{i}(\tilde{\pi}^{i}\times\pi^{-i};s)-\epsilon$ (2)
for all $\tilde{\pi}^{i}:S\rightarrow\Delta(A^{i})$, $i\in[N]$ and $s\in S$ (Daskalakis et al., 2023, Definition 2.1).
We also say that $\epsilon$-CCE is $\epsilon$-approximate perfect Markov Nash equilibrium if the joint strategy is also a product strategy, i.e., $\pi=\bigtimes_{i}\pi^{i}$ for some $\pi^{i}:S\rightarrow\Delta(A^{i})$ and $i\in[N]$ (Daskalakis et al., 2023, Definition 2.2). In stochastic games, there always exists perfect Markov Nash equilibrium (also known as Markov perfect equilibrium) (Fink, 1964), and thus so do the perfect Markov coarse correlated equilibrium. However, there might be multiple equilibria leading to different outcomes with respect to certain performance metrics.
Definition 2.
We say that a joint Markov stationary strategy $\pi:S\rightarrow\Delta(A)$ is $\epsilon$-efficient with respect to $U(\cdot)$ if $U(\pi;s)\geq U(\tilde{\pi};s)-\epsilon$ for all $\tilde{\pi}:S\rightarrow\Delta(A)$ and $s\in S$.
In identical-interest stochastic games (also known as stochastic teams) where $r^{i}(s,a)=r(s,a)$ for all $(s,a)$ for some common $r:S\times A\rightarrow\mathbb{R}$, $\epsilon$-CCE may lead to arbitrarily poor performance with respect to the agents’ common objective $U(\cdot)=U^{i}(\cdot)$ for all $i$, compared to the solutions where agents are controlled centrally to maximize the common objective $U(\cdot)$. The following proposition shows the existence of efficient equilibrium in stochastic teams, and its proof follows from Definitions 1 and 2.
Proposition 1.
Given a stochastic team characterized by $\langle S,(A^{i})_{i\in[N]},r,p,\gamma\rangle$ with the common objective $U(\cdot)$. A joint Markov stationary strategy $\pi$ that is $\epsilon$-efficient with respect to $U(\cdot)$ is also $\epsilon$-CCE of the underlying game.
In this paper, we present learning dynamics reaching such efficient equilibria. To this end, we focus on the stage-game framework, going back to the introduction of stochastic games by (Shapley, 1953). Particularly, given the joint Markov stationary strategy $\pi$, let $v_{\pi}^{i}:S\rightarrow\mathbb{R}$ be the agent $i$’s value function such that $v^{i}_{\pi}(s)$ is the value of state $s$ for agent $i$ if the game starts at state $s$ and the agents play according to $\pi$, i.e., $v^{i}_{\pi}(s)=U^{i}(\pi;s)$. Correspondingly, let $Q^{i}_{\pi}:S\times A\rightarrow\mathbb{R}$ be the agent $i$’s Q-function such that $Q^{i}_{\pi}(s,a)$ is the value of state-action pair $(s,a)$ for agent $i$. By (1), the Bellman operator for a strategy yields that
$Q^{i}_{\pi}(s,a)=r^{i}(s,a)+\gamma\sum_{s^{\prime}\in S}p(s^{\prime}\mid s,a)\cdot v^{i}_{\pi}(s^{\prime})\quad\forall(s,a),$ (3a)
$v^{i}_{\pi}(s)=\mathrm{E}_{a\sim\pi(s)}\left[Q^{i}_{\pi}(s,a)\right]\quad\forall s.$ (3b)
Then, we can view stochastic games as the repeated play of stage games whose payoffs are the Q-functions associated with how they play in these stage games, whenever the associated state gets visited.
3 Logit-Q Dynamics for Stochastic Games
In this section, we introduce the logit-Q dynamics within the stage-game framework. In this framework, agents follow (independent) log-linear learning in each stage game as if the stage-game payoffs are the Q-functions estimated. They estimate the Q-functions associated with how they play in these stage games by solving the fixed-point problem (3) iteratively based on the game history. To this end, each agent $i$ keeps track of
  • $a_{t}^{j}(s)\in A^{j}$ for each $j\in[N]$ denoting agent $j$’s latest action at state $s$ until and including stage $t$, with a slight abuse of notation,
  • $c_{t}(s,a)$ and $c_{t}(s)\in\{0,1,\ldots\}$ denoting, resp., the number of times the pair $(s,a)$ and $s$ get realized until and including stage $t$,
  • $Q_{t}^{i}(s,a)\in\mathbb{R}$ denoting the Q-function estimate of the pair $(s,a)$ at stage $t$.
Recall the friction in the action updates of the log-linear learning, also known as the lock-in property. To model such friction, we consider an i.i.d. process $\{I_{t}\subset A\}_{t\geq 0}$ determining the agents updating their actions. For example, agent $i$ revises his/her action at stage $t$ if and only if $i\in I_{t}$. For the distribution of the revision process, we consider two schemes:
  • In the classical coordinated scheme, we have $I_{t}\sim\mathcal{P}^{\mathrm{coor}}$ and [^2] For a finite set $I$, we denote its number of elements by $|I|$.
    $\mathcal{P}^{\mathrm{coor}}(I):=\mathds{1}_{\{|I|=1\}}\frac{1}{N}\quad\forall I\subset[N],$ (4)
    i.e., only a randomly selected agent can update his/her action.
  • In the independent scheme, we have $I_{t}\sim\mathcal{P}^{\mathrm{ind}}$ and
    $\mathcal{P}^{\mathrm{ind}}(I)=\delta^{|I|}(1-\delta)^{N-|I|}\quad\forall I\subset[N],$ (5)
    i.e., agents update their actions independently with probability $\delta\in(0,1)$.
Let $s_{t}\in S$ be the state at stage $t$. Each agent $i\in I_{t}$ revises his/her action according to the smoothed best response as if the game payoffs are $Q_{t}^{i}(s_{t},\cdot)$ and the other agents are playing the latest actions $a_{t-1}^{-i}(s_{t})=\{a_{t-1}^{j}(s_{t})\}_{j\neq i}$. Particularly, given $q\in\mathbb{R}^{B}$ for some finite set $B$, we define the smoothed best response by
$\underline{\mathrm{br}}(q)(b):=\frac{\exp(q(b)/\tau)}{\sum_{b^{\prime}\in B}\exp(q(b^{\prime})/\tau)}\quad\forall b\in B.$ (6)
Then, we have $a_{t}^{i}\sim\underline{\mathrm{br}}(Q_{t}^{i}(s_{t},\cdot,a_{t-1}^{-i}(s_{t})))$ and the parameter $\tau>0$ balances exploration vs exploitation. For example, agent $i$ plays best response(s) with higher probabilities (whose sum goes to $1$ as $\tau\rightarrow 0$) and can explore any action with some positive probabilities (going to $1/|A^{i}|$ as $\tau\rightarrow\infty$). On the other hand, each agent $j\notin I_{t}$, not revising their actions, plays their latest action at that state, i.e., $a_{t}^{j}=a_{t-1}^{j}(s_{t})$.
Agents estimate the joint Markov stationary strategy $\pi:S\rightarrow\Delta(A)$, corresponding to how they play in these stage games based on the game history. To this end, we consider two alternatives:
  • In the averaging approach, agents take the empirical average of the joint actions played in each state visited. Then, the estimated play is given by
    $\pi_{t}^{\mathrm{ave}}(s)(a)=\frac{c_{t}(s,a)}{c_{t}(s)}\quad\forall(s,a).$ (7)
  • In the exploration-free approach, agents focus on the joint actions that get played the most frequently to mitigate the impact of exploration inherent to the smoothed best response (due to $\tau>0$ in (6)) as the best responses get played with higher probabilities. Then, the estimated play is given by
    $\pi_{t}^{\mathrm{free}}(s)(a)=\left\{\begin{array}[]{ll}1/|A_{t}(s)|&\mbox{if }a\in A_{t}(s)\\ 0&\mbox{if }a\notin A_{t}(s)\end{array}\right.$ (8)
    for all $(s,a)$, where the set $A_{t}(s):=\operatorname*{argmax}_{a}\{c_{t}(s,a)\}\subset A$.
Let $\pi_{t}^{\mathrm{ave}}(s)$ and $\pi_{t}^{\mathrm{free}}(s)$ be uniform distributions for $c_{t}(s)=0$, without loss of generality.
To compute the Q-function associated with the estimated play, the agents can solve (3) iteratively and use one-stage look ahead to address the unknown transition kernel, as in the classical Q-learning. After observing the next state $s_{t+1}$ at stage $t+1$, each agent $i$ updates his/her Q-function estimate for the pair $(s_{t},a_{t})$ according to
$Q_{t+1}^{i}(s_{t},a_{t})=Q_{t}^{i}(s_{t},a_{t})+\beta_{c_{t}(s_{t},a_{t})}\left(r^{i}_{t}+\gamma v_{t}^{i}(s_{t+1})-Q_{t}^{i}(s_{t},a_{t})\right),$ (9)
where $r_{t}^{i}:=r^{i}(s_{t},a_{t})$ is the stage-payoff received and $v_{t}^{i}$ denotes the value function estimate. Here, the agents use the value $v_{t}^{i}(s_{t+1})$ of the next state as an approximation of the expected continuation payoff $\sum p(s^{\prime}\mid s,a)v_{t}^{i}(s^{\prime})$ in (3a). Given the estimated play $\pi_{t}$, the value $v_{t}^{i}$ is given by
$v_{t}^{i}(s_{t+1})=\mathrm{E}_{a\sim\pi_{t}(s_{t+1})}[Q_{t}^{i}(s_{t+1},a)]$ (10)
due to (3b). The reference step size $\beta_{c}\in(0,1)$ for all $c=0,1,\ldots$ implies that the agents update their Q-function estimates to convex combinations of the previous estimates $Q_{t}^{i}$ and the stage-games’ (estimated) payoffs $r^{i}_{t}+\gamma\mathrm{E}_{a\sim\pi_{t}(s_{t+1})}[Q_{t}^{i}(s_{t+1},a)]$. Since agents do not observe the counterfactuals about the joint actions that are not played and the states that are not visited in the current stage, they do not update the other Q-function estimates, i.e., $Q_{t+1}^{i}(s,a)=Q_{t}^{i}(s,a)$ for all $(s,a)\neq(s_{t},a_{t})$.
Algorithm 1 Logit-Q Dynamics
0:  each agent can observe $(s_{t},a_{t})$ and keep track of $a_{t}(\cdot)$ and $c_{t}(\cdot)$
1:  Input: initializations $Q^{i}_{0}(s,a)=0$ for all $(i,s,a)\in[N]\times S\times A$.
2:  for each stage $t=0,1,\ldots$ do
3:     $s_{t}$ gets realized
4:     if $t=0$ then jump to Step 6
5:     each agent $i$ updates the Q-function estimates for stage $t-1$ according to
$Q_{t}^{i}(s,a)=Q_{t-1}^{i}(s,a)+\mathds{1}_{\{(s,a)=(s_{t-1},a_{t-1})\}}\cdot\beta_{c_{t-1}(s,a)}\cdot(\widehat{Q}_{t}^{i}-Q_{t-1}^{i}(s,a)),$
where $\widehat{Q}_{t}^{i}:=r^{i}_{t-1}+\gamma\mathrm{E}_{a\sim\pi_{t-1}(s_{t})}[Q_{t-1}^{i}(s_{t},a)]$ and $\pi_{t-1}(\cdot)$ depends on $c_{t-1}(\cdot)$
6:     $I_{t}\sim\mathcal{P}$ gets realized
7:     each agent $i\in I_{t}$ plays $a^{i}_{t}\sim\underline{\mathrm{br}}(Q_{t}^{i}(s_{t},\cdot,a_{t-1}^{-i}(s_{t})))$
8:     each agent $i\notin I_{t}$ plays $a^{i}_{t}=a_{t-1}^{i}(s_{t})$
9:  end for
Table 1: A family of the logit-Q dynamics with the parameters $\mathcal{P}$ and $\pi_{t}$, as described in (4), (5), (7),(8)
Logit-Q Dynamics Revision Process $\mathcal{P}$ Estimated Play $\pi_{t}$
Averaged Logit-Q (AL-Q) $\mathcal{P}^{\mathrm{coor}}$ $\pi_{t}^{\mathrm{ave}}$
Exploration-free Logit-Q (EL-Q) $\mathcal{P}^{\mathrm{coor}}$ $\pi_{t}^{\mathrm{free}}$
Averaged Independent Logit-Q (AIL-Q) $\mathcal{P}^{\mathrm{ind}}$ $\pi_{t}^{\mathrm{ave}}$
Exploration-free Independent Logit-Q (EIL-Q) $\mathcal{P}^{\mathrm{ind}}$ $\pi_{t}^{\mathrm{free}}$
Algorithm 1 explicitly describes a family of logit-Q dynamics, varying in terms of the revision processes $\mathcal{P}$ and the estimated plays $\pi_{t}$, as tabulated in Table 1. Recall that agents update their Q-function estimates after observing the next state at the next stage. We also consider that agents initialize their Q-function estimates as $Q_{0}^{i}(s,a)=0$ for all $(i,s,a)$, rather than choosing arbitrarily. Such coordination can be a challenge for ad hoc team formations. However, the step sizes $\beta_{c}\in(0,1)$ suppress the impact of the initialization on the estimates at the future stages and we examine the impact of arbitrary initializations numerically later in Section 6. Furthermore, such initialization and $\beta_{c}\in(0,1)$ for all $c$ yield that the Q-function estimates are always bounded by $0\leq Q_{t}^{i}(s,a)\leq 1/(1-\gamma)$ for all $(i,s,a)$ since $r^{i}:S\times A\rightarrow[0,1]$. Correspondingly, agents revising their actions can take any action with some positive probability uniformly bounded away from zero by (6).
4 Convergence Results
In this section, we characterize the convergence properties of Algorithm 1 based on the following assumption on the step size $\beta_{c}$.
Assumption 1.
The reference step size $\beta_{c}\in(0,1)$ satisfies
  1. $(i)$
    Monotonic decay: $\beta_{c}\rightarrow 0$ as $c\rightarrow\infty$ and $\beta_{c}>\beta_{c^{\prime}}$ for all $c<c^{\prime}$,
  2. $(ii)$
    Sufficiently slow decay: $\sum_{c}\beta_{c}=\infty$,
  3. $(iii)$
    Sufficiently fast decay: $c\beta_{\lfloor mc\rfloor}\rightarrow 0$ as $c\rightarrow\infty$ for any $m\in(0,1]$. [^3] Let $\lfloor r\rfloor:=\{z\in\mathbb{Z}:z\leq r\}$ for any $r\in\mathbb{R}$ denote the floor function.
Assumption 1-$(i)$ and Assumption 1-$(ii)$ are standard in stochastic approximation, and Assumption 1-$(iii)$ is standard for asynchronous two-timescale stochastic approximation if we view $1/(c+1)$ as the step size for the dynamics evolving at the fast timescale (Perkins and Leslie, 2012, Section 4.2). For example, $\beta_{c}=1/((c+1)\log(c+1))$ satisfies Assumption 1. We also note that Assumption 1-$(iii)$ implies $\sum_{c}\beta_{c}^{2}<\infty$, playing an important role in addressing the stochastic approximation noise.
The following theorem characterizes the convergence properties of the logit-Q dynamics for stochastic teams.
Table 2: Error bounds in Theorem 1, where $\Lambda(\delta)$ and $\Xi(\delta)$ as described later in Proposition 3 and (54), respectively. Both $\Lambda(\delta)$ and $\Xi(\delta)$ decay to zero as $\delta\rightarrow 0^{+}$.
AL-Q AIL-Q EL-Q EIL-Q
$e_{Q}$ in (11a) $\frac{\gamma}{1-\gamma}\cdot(\tau\log|A|)$ $\frac{\gamma}{1-\gamma}\cdot(\tau\log|A|+\frac{1}{1-\gamma}\cdot\Lambda(\delta))$ 00 $\frac{\gamma}{1-\gamma}\cdot\Xi(\delta)$
Theorem 1.
Given a stochastic team $\langle S,(A^{i})_{i\in[N]},r,p,\gamma\rangle$ with the common objective $U(\cdot)$, consider that every agent follow Algorithm 1 with the same revision-process and estimated-play schemes, as in Table 1. Suppose that Assumption 1 holds and the underlying stochastic game is irreducible in the sense that given any pure Markov stationary strategy, the states visited form an irreducible Markov chain. [^4] Such irreducibility assumptions are widely used for stochastic games, e.g., see (Hoffman and Karp, 1966; Brafman and Tennenholtz, 2002; Sayin et al., 2021). Then, for each $i\in[N]$, we have
$\limsup_{t\rightarrow\infty}\left|\,Q_{t}^{i}(s,a)-Q_{*}(s,a)\,\right|\leq e_{Q}\quad\forall(s,a)$ (11a)
$\limsup_{t\rightarrow\infty}\left|\max_{\pi}\{U(\pi;s)\}-U(\pi_{t};s)\,\right|\leq\frac{1+\gamma}{\gamma(1-\gamma)}\cdot e_{Q}\quad\forall s$ (11b)
almost surely, where the team-optimal Q-function $Q_{*}:S\times A\rightarrow\mathbb{R}$ is the unique fixed point of the following Q-function variant of the Bellman optimality operator:
$Q_{*}(s,a)=r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}\mid s,a)\max_{a^{\prime}\in A}\;\{Q_{*}(s^{\prime},a^{\prime})\}\quad\forall(s,a)$ (12)
and the error bound $e_{Q}\geq 0$ is as tabulated in Table 2 for different revision-process and estimated-play schemes.
Theorem 1 shows that in Algorithm 1, the Q-function estimates and the estimated plays almost surely converge, resp., to the team-optimal Q-functions and efficient stationary equilibrium approximately, with quantifiable error bounds (see Table 2). The proof is moved to Section 5.
The following corollary to Theorem 1 shows the rationality of Algorithm 1 against others playing according to pure stationary strategies.
Corollary 1.
Given a general-sum stochastic game $\langle S,(A^{i},r^{i})_{i\in[N]},p,\gamma\rangle$, consider that agent $i$ follows AIL-Q or EIL-Q while the other agents $j\neq i$ play according to pure Markov stationary strategy, e.g., $\pi^{-i}:S\rightarrow A^{-i}$. Suppose that Assumption 1 holds and the states visited under the strategy $\tilde{\pi}^{i}\times\pi^{-i}$ form an irreducible Markov chain for any pure Markov stationary strategy $\tilde{\pi}^{i}:S\rightarrow A^{i}$. Then, we have
$\limsup_{t\rightarrow\infty}|Q_{t}^{i}(s,a^{i},\pi^{-i}(s))-q_{*}^{i}(s,a^{i})|\leq e_{Q}\quad\forall(s,a^{i})$ (13a)
$\limsup_{t\rightarrow\infty}|\max_{\tilde{\pi}^{i}}\{U^{i}(\tilde{\pi}^{i}\times\pi^{-i};s)\}-U^{i}(\pi_{t};s)|\leq\frac{1+\gamma}{\gamma(1-\gamma)}\cdot e_{Q}\quad\forall s$ (13b)
almost surely, where $q_{*}^{i}:S\times A^{i}\rightarrow\mathbb{R}$ is the local Q-function associated with the best response to $\pi^{-i}$ and is the unique fixed point of the Bellman operator variant
$q_{*}^{i}(s,a^{i})=r^{i}(s,a^{i},\pi^{-i}(s))+\gamma\sum_{s^{\prime}}p(s^{\prime}|s,a^{i},\pi^{-i}(s))\max_{\tilde{a}^{i}\in A^{i}}\{q_{*}^{i}(s^{\prime},\tilde{a}^{i})\}\quad\forall(s,a^{i}).$ (14)
The error bound $e_{Q}\geq 0$ is as in Table 2 for $A=A^{i}$.
Proof.
We can consider the underlying stochastic game as a single-agent stochastic team characterized by $\langle S,A^{i},\bar{r}^{i},\bar{p},\gamma\rangle$ where $\bar{r}^{i}(s,a^{i})=r^{i}(s,a^{i},\pi^{-i}(s))$ and $\bar{p}(s^{\prime}|s,a^{i})=p(s^{\prime}|s,a^{i},\pi^{-i}(s))$. Then, the proof follows from Theorem 1. ∎
4.1 Beyond Stochastic Teams
Next, consider an $N$-agent game characterized by $\mathcal{G}=\langle A^{i},u^{i}\rangle_{i\in[N]}$, where $A^{i}$ is the finite action set of agent $i$, $A:=\prod_{j\in[N]}A^{j}$ is the set of action profiles, and $u^{i}:A\rightarrow\mathbb{R}$ is the payoff function of agent $i$. We say that $\mathcal{G}$ is a potential game if there exists a potential function $\Phi:A\rightarrow\mathbb{R}$ satisfying
$\Phi(a^{i},a^{-i})-\Phi(\tilde{a}^{i},a^{-i})=u^{i}(a^{i},a^{-i})-u^{i}(\tilde{a}^{i},a^{-i})\quad\forall(a^{i},a^{-i},\tilde{a}^{i})\mbox{ and }i\in[N].$ (15)
Note that identical-interest games are potential games with the common payoff as the potential function.
For potential games, if agents follow (independent) log-linear learning dynamics, then by (6) and (15), the smoothed best response can be written as
$\frac{\exp(u^{i}(a^{i},a^{-i})/\tau)}{\sum_{\tilde{a}^{i}\in A^{i}}\exp(u^{i}(\tilde{a}^{i},a^{-i})/\tau)}=\frac{\exp(\Phi(a^{i},a^{-i})/\tau)}{\sum_{\tilde{a}^{i}\in A^{i}}\exp(\Phi(\tilde{a}^{i},a^{-i})/\tau)}.$ (16)
This implies that we can view the (independent) log-linear learning dynamics as if the agents have the potential function as their objective without loss of generality. Therefore, the convergence of the (independent) log-linear learning dynamics to near efficient equilibrium with respect to the potential function follows from the near efficient equilibrium convergence of the dynamics in identical-interest games.
However, the convergence of Algorithm 1 to near efficient stationary equilibrium in stochastic teams, as shown in Theorem 1, does not imply the convergence in cases where the stage-payoffs induce a potential game in general. We can list the challenges as follows: $(i)$ the stage games may not be potential games even though the stage-payoffs induce potential games and $(ii)$ even when stage games are potential games, reaching the efficient equilibrium in the stage games with respect to the potential function does not imply the maximization of the local stage-game payoffs (which plays an important role for the contraction property).
The following corollary to Theorem 1 (similar to (Sayin et al., 2022b, Corollary 4)) shows that we can address these challenges for a certain class of stochastic games.
Corollary 2.
Consider a general-sum stochastic game $\langle S,(A^{j},r^{j})_{j\in[N]},p,\gamma\rangle$ where for some specific agent $i$, the stage-payoffs of agents $j\neq i$ satisfy
$r^{i}(s,a^{j},a^{-j})-r^{i}(s,\tilde{a}^{j},a^{-j})=r^{j}(s,a^{j},a^{-j})-r^{j}(s,\tilde{a}^{j},a^{-j})\quad\forall(s,a^{j},a^{-j},\tilde{a}^{j})$ (17)
similar to (15), and the underlying transition kernel satisfies
$p(s^{\prime}\mid s,a^{i},a^{-i})=p(s^{\prime}\mid s,a^{i},\tilde{a}^{-i})\quad\forall(s,a^{i},a^{-i},\tilde{a}^{-i},s^{\prime})$ (18)
such that only agent $i$ controls the state transitions.
Suppose that every agent follow Algorithm 1, Assumption 1 holds and the underlying stochastic game is irreducible, as described in Theorem 1. Then, for the single-controller agent $i$, we have
$\limsup_{t\rightarrow\infty}\left|\;Q_{t}^{i}(s,a)-Q_{*}^{i}(s,a)\;\right|\leq e_{Q}\quad\forall(s,a)$ (19a)
$\limsup_{t\rightarrow\infty}\left|\;\max_{\pi}\{U^{i}(\pi;s)\}-U^{i}(\pi_{t};s)\;\right|\leq\frac{1+\gamma}{\gamma(1-\gamma)}\cdot e_{Q}\quad\forall s$ (19b)
almost surely, where the Q-function $Q_{*}^{i}:S\times A\rightarrow\mathbb{R}$ is the unique fixed point of the following Bellman operator variant:
$Q_{*}^{i}(s,a)=r^{i}(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}\mid s,a)\max_{a^{\prime}\in A}\;\{Q_{*}^{i}(s^{\prime},a^{\prime})\}\quad\forall(s,a).$ (20)
The error bounds are as described in Table 2.
Proof.
By (17), the stage-payoff functions $\{r^{j}(s,\cdot)\}_{j\in[N]}$ induce a potential game whose potential function is $r^{i}(s,\cdot)$ for each $s$. Then, the Q-update (9) yields that stage-game payoffs $\{Q_{t}^{j}(s,\cdot)\}_{j\in[N]}$ induce a potential game whose potential function is $Q_{t}^{i}(s,\cdot)$ since only agent $i$ can control the continuation payoff due to the single-controller condition (18). Therefore, the proof follows from Theorem 1 since we can view the underlying stochastic game as the stochastic team $\langle S,(A^{j})_{j\in[N]},r^{i},p,\gamma\rangle$. ∎
5 Proof of Theorem 1
The proof follows from showing that the estimated plays $\pi_{t}$’s track the (soft) maximizers of the Q-function estimates $Q_{t}^{i}$’s. Particularly, if the estimated plays track the (soft) maximizers, then the Q-update (9) approximates the (soft) Q-learning algorithm, where the entire team coordinates in the (soft) maximizer, and the Q-learning is known to converge to the Q-function associated with the optimal team strategy under standard conditions on step sizes and transition probabilities.
In the following, we show this tracking result by focusing on the evolution of the estimated plays across the repeated play of stage games. Recall that in the logit-Q dynamics, agents follow the logit dynamics in the stage games specific to each state whenever the associated state gets visited as if the stage-game payoffs are their Q-function estimates. Hence, we first characterize the convergence properties of the logit dynamics for non-stationary environments, which might be of independent interest.
5.1 Convergence Properties of the Logit Dynamics
In the logit dynamics, agents revise their actions according to a revision process, e.g., either in a coordinated or independent way. The agents revising their actions play according to the smoothed best response to the others’ previous actions.
Notably, the action profiles played in the logit dynamics form an irreducible and aperiodic Markov chain (Marden and Shamma, 2012). The following proposition characterizes the dependence of the (long-run) behavior of this Markov chain on the game payoffs.
Proposition 2.
Consider that agents follow the logit dynamics in the repeated play of a potential game $\mathcal{G}=\langle(A^{i},u^{i})_{i\in[N]}\rangle$ with the potential function $\Phi$. Then, the transition matrix and the unique stationary distribution for the Markov chain formed by the action profiles are locally Lipschitz continuous functions of the potential function.
Proof.
In the logit dynamics, the transition probabilities of the Markov chain formed by the action profiles can be written as
$\Pr(a_{+}\mid a)=\sum_{I\subset[N]}\mathcal{P}(I)\cdot\prod_{i\in I}\underline{\mathrm{br}}(\Phi(\cdot,a^{-i}))(a_{+}^{i})\cdot\prod_{j\notin I}\mathds{1}_{\{a_{+}^{j}=a^{j}\}}\quad\forall(a,a_{+})\in A\times A,$ (21)
by (16), where $\mathcal{P}$ is the distribution of the revision process, e.g., as described in (4) or (5). Since $\underline{\mathrm{br}}(\cdot)$ is a continuously differentiable function, (21) yields that the transition matrix is a locally Lipschitz continuous function of the potential function. Furthermore, in irreducible and aperiodic Markov chains, the unique stationary distributions are Lipschitz continuous functions of their transition matrices (Cho and Meyer, 2001, Section 3). ∎
This proposition plays an important role in the convergence analysis since the stage-game payoffs, i.e., Q-function estimates, can evolve in time due to (9). Small changes in the Q-function estimates can result in (relatively) small changes in the transition probabilities and the corresponding stationary distributions due to the Lipschitz continuity.
Figure 1: Consider two states $s$ and $\overline{s}$ with action profiles $\{a,b\}$ and $\{\overline{a},\overline{b}\}$, respectively. This is a figurative illustration for the global trajectory of the pairs $(s_{t},a_{t})_{t\geq 0}$ evolving over the global timescale $t=0,1,\ldots$, and the local trajectory of the pairs $(w_{k},z_{k})_{k\geq 0}$ specific to $s$ and evolving over the local timescale $k=1,2,\ldots$ whenever $s$ gets visited. Let $t_{k}$ be the time of the $k$th visit to $s$. In the local trajectory, $w_{k}$ and $z_{k}$ denote, resp., the action profile played at the $k$th visit to $s$ and the trajectory of the state-action pairs realized in between the $k$th and $(k+1)$th visit to $s$.
In the following, we quantify the impact of these changes on the long-run behavior of the dynamics in comparison to the cases where the underlying stage game is stationary. However, the coupling of the dynamics across states poses a challenge. To address this challenge, we focus on the dynamics specific to a fixed state $s$ and model the dynamics specific to other states as some external (possibly coupled) processes. Let $w_{k}$ and $z_{k}$ denote, resp., the action profile played at the $k$th visit to $s$ and the trajectory of the state-action pairs realized in between the $k$th and $(k+1)$th visit to $s$, as illustrated in Figure 1. Then, we have $w_{k}\in A$ and $z_{k}\in Z$, where $Z$ represents the countable set of all possible trajectories in between consecutive visits to $s$.
Let $t_{k}$ be the stage for the $k$th visit to $s$. Then, we have $\sigma\left((s_{\tau},a_{\tau})_{\tau<t_{k}},s_{t_{k}}\right)=\sigma\left((w_{l},z_{l})_{l<k}\right)$. The Q-function estimate $Q_{t_{k}}$ is $\sigma\left((w_{l},z_{l})_{l<k}\right)$-measurable since $Q_{t_{k}}$ is $\sigma\left((s_{\tau},a_{\tau})_{\tau<t_{k}},s_{t_{k}}\right)$-measurable by (9). Therefore, the randomness on $(w_{k},z_{k})$ is contingent only on $(w_{l},z_{l})_{l<k}$ and there exists a transition kernel $\mathbb{P}$ such that $(w_{k},z_{k})\sim\mathbb{P}\left(\cdot\mid(w_{l},z_{l})_{l<k}\right)$. The following lemma can characterize the long-run behavior of the logit dynamics in such non-stationary environments by coupling its evolution with the logit dynamics in stationary environments.
Lemma 1.
Consider two discrete-time processes $\{\widehat{w}_{k}\}_{k\geq 0}$ and $\{(w_{k},z_{k})\}_{k\geq 0}$ over a finite set $W$ and the product space $W\times Z$ with some countable space $Z$. The processes $\{\widehat{w}_{k}\}$ and $\{(w_{k},z_{k})\}$ evolve, resp., according to $\widehat{w}_{k}\sim\widehat{\mathbb{P}}(\cdot|\widehat{w}_{k-1})$ and $(w_{k},z_{k})\sim\mathbb{P}(\cdot|(w_{l},z_{l})_{l<k})$. Let $\mathbb{P}^{w}(\cdot|(w_{l},z_{l})_{l<k})$ denote the marginal of $\mathbb{P}(\cdot|(w_{l},z_{l})_{l<k})$ over $Z$ such that $w_{k}\sim\mathbb{P}^{w}(\cdot|(w_{l},z_{l})_{l<k})$. Suppose that these processes satisfy the following conditions:
Condition $(i)$ There exist $\widehat{\varepsilon},\varepsilon>0$, a destination $w^{\prime}\in W$ and $\kappa\in\mathbb{N}$ such that given any $(\widehat{w},w,z)\in W\times W\times Z$, where $\widehat{w}\neq w$, the processes $\{\widehat{w}_{k}\}$ and $\{(w_{k},z_{k})\}$ can follow, resp., $\kappa$-length paths $\widehat{w}=\widehat{h}^{0},\dots,\widehat{h}^{\kappa}=w^{\prime}$, and $(w,z)=(h^{0},g^{0}),\dots,(h^{\kappa},g^{\kappa})=(w^{\prime},z^{\prime})$ for some $z^{\prime}$, where $\widehat{h}^{l}\neq h^{l}$ for all $l<\kappa$. Furthermore, for all $k$, we have
$\Pr\left(\bigwedge_{l=1}^{\kappa}\widehat{w}_{k+l}=\widehat{h}^{l}\mid\widehat{w}_{k}=\widehat{w}\right)\geq\hat{\varepsilon},\Pr\left(\bigwedge_{l=1}^{\kappa}w_{k+l}=h^{l}\mid(w_{k},z_{k})=(w,z)\right)\geq\varepsilon.$ (22)
Condition $(ii)$ For any $k=0,\ldots,K-1$, there exists $\lambda\in(0,1]$ such that
$\left\|\mathbb{P}^{w}\left(\cdot\mid(w_{l},z_{l})_{l\leq k}\right)-\widehat{\mathbb{P}}(\cdot\mid w_{k})\right\|_{\mathrm{TV}}\leq 1-\lambda\quad\forall(w_{l},z_{l})_{l\leq k}.$ (23)
Let $\widehat{\mu}_{k}\in\Delta(W)$ and $\mu_{k}\in\Delta(W\times\mathbb{Z})$ denote, resp., the distributions of $\widehat{w}_{k}$ and $(w_{k},z_{k})$. Furthermore, $\widehat{\mu}^{w}\in\Delta(W)$ denotes the marginal of $\mu_{k}$ over $Z$. Then, there exists a coupling over $\{\widehat{w}_{k}\}$ and $\{(w_{k},z_{k})\}$ such that
$\|\widehat{\mu}_{k}-\mu_{k}^{w}\|_{1}\leq 2(1-\varepsilon\widehat{\varepsilon})^{\frac{k}{\kappa}-1}+2(1-\lambda^{\kappa})\frac{1+\varepsilon\widehat{\varepsilon}}{\varepsilon\widehat{\varepsilon}}\quad\forall t=0,\ldots,K-1.$ (24)
We highlight that the first term on the right-hand side of (24) goes to zero as $k\rightarrow\infty$ whereas the second term is a constant depending on the distance between the transition probabilities. For example, if $\lambda=1$ for $K\rightarrow\infty$, then the processes $\{\widehat{w}_{k}\}$ and $\{w_{k}\}$ are identical Markov chains and the second term is zero. Correspondingly, the proof of Lemma 1 (provided in Appendix A) has a flavor similar to the proof of the Ergodicity Theorem for Markov chains (Levin and Peres, 2017, Chapter 4).
Remark 1.
The literature on perturbation bounds for the stationary distribution of Markov chains are mainly for homogeneous chains and based on the maximum entry-wise perturbation to address general cases, e.g., see (Cho and Meyer, 2001; Liu, 2012). However, these bounds are not tight enough for our analysis. Hence, in Lemma 1, we focus on the structure of these chains for sharper perturbation bounds while also addressing the non-homogeneous cases through finite-time analysis. Therefore, Lemma 1 can be of independent interest.
Lastly, the following proposition characterizes the stationary distributions of the action profiles in the classical and independent learning dynamics based on Lemma 1. A preliminary version of this lemma appeared in (Unlu and Sayin, 2023).
Proposition 3.
Consider that agents follow the logit dynamics in the repeated play of a potential game $\mathcal{G}=\langle(A^{i},u^{i})_{i\in[N]}\rangle$ with the potential function $\Phi$. Let $\mu^{\mathrm{coor}},\mu^{\mathrm{ind}}\in\Delta(A)$ denote the unique stationary distributions of the Markov chains formed by the actions profiles for the coordinated and independent revision processes, as described in (4) and (5), respectively. Then, we have $\mu^{\mathrm{coor}}=\underline{\mathrm{br}}(\Phi(\cdot))$ and
$\|\mu^{\mathrm{coor}}-\mu^{\mathrm{ind}}\|_{1}\leq\Lambda(\delta):=2(1-\lambda(\delta))\cdot\frac{1+\lambda(\delta)\cdot\epsilon^{2}}{\lambda(\delta)\cdot\epsilon^{2}}.$ (25)
For independent-revision schemes, $\lambda(\delta)\in(0,1)$ is the probability that only one agent revises actions for $N$ stages conditioned on that at least one agent revises actions and $\epsilon\in(0,1)$ is a uniform lower bound on the probability for any $N$-length trajectory where only one agent revises actions, which depends on the inherent exploration.
As $\delta\rightarrow 0^{+}$, we have $\Lambda(\delta)\rightarrow 0$ since $\lambda(\delta)\rightarrow 1$. We also highlight the resemblance of (25) to the second term on the right-hand side of (24). Correspondingly, the proof of Proposition 3 (provided in Appendix D) follows from Lemma 1 through the quantification of the distance between the transition matrices of the coordinated and independent-revision schemes.
In the following, we leverage Propositions 2 and 3 and Lemma 1 to prove that the estimated plays can track the (soft) maximizers of the Q-function estimates.
5.2 Tracking Result for the Estimated Plays
For the estimated plays, we can focus on $\pi^{\mathrm{ave}}_{t}$ since we can compute $\pi^{\mathrm{free}}_{t}$ based on $\pi^{\mathrm{ave}}_{t}$ according to
$\pi^{\mathrm{free}}_{t}(s)(a)>0\quad\Leftrightarrow\quad a\in A_{t}(s)=\operatorname*{argmax}_{\tilde{a}\in A}\{\pi^{\mathrm{ave}}_{t}(s)(a)\}\quad\forall(s,a).$ (26)
Let $\mu(u)\in\Delta(A)$ denote the unique stationary distribution of the action profiles for the logit dynamics followed in the identical-interest game $\mathcal{G}=\langle(A^{i})_{i\in[N]},u\rangle$. Then, the following proposition shows that the empirical averages $\pi^{\mathrm{ave}}_{t}$ can track the stationary distributions associated with the stage-game payoffs $Q_{t}$’s.
Proposition 4 (Tracking Result).
We have $\lim_{t\rightarrow\infty}\|\pi^{\mathrm{ave}}_{t}(s)-\mu(Q_{t}(s,\cdot))\|_{1}=0$ almost surely for each $s$.
Proof.
If the stage-game payoffs, i.e., the Q-function estimates, were stationary, the empirical average $\pi^{\mathrm{ave}}_{t}$ would have converged to the stationary distribution of the homogeneous Markov chain formed by the action profiles due to the ergodicity of the underlying chain. Even though the Q-function estimates evolve in time, the vanishing step size by Assumption 1-$(i)$ and the boundedness of the iterates by the nature of the update (9) yield that one-step changes in the Q-function estimates decay to zero. The changes in the transition probabilities also remain (relatively) small for small changes in the game payoffs in the logit dynamics by Proposition 2. Therefore, the stage games become quasi-stationary within certain time periods where the accumulated changes in the transition probabilities are small. However, longer periods result in larger accumulated changes. We can address these seemingly conflicting conditions by dividing the horizon into epochs whose lengths grow at a sufficiently slow rate.
More explicitly, fix some state $s$ and divide the horizon into epochs such that the state $s$ gets visited only for a certain number of times specific to the epoch and the first state of every epoch is $s$. For clear exposition, we use $k$ and $(n)$, resp., as subscripts for the parameters associated with the $k$th visit to $s$ and the $n$th epoch for $s$. For example, $t_{k}$, $t_{(n)}$, and $k_{(n)}$ denote the stage that $s$ gets visited for the $k$th time, the stage that the $n$th epoch starts, and the number of times $s$ gets visited until and including stage $t_{(n)}$, respectively.
For each epoch $n$, we define the filtration $\mathcal{F}_{(n)}=\sigma(\{s_{t},a_{t}\}_{t<t_{(n)}},s_{t_{(n)}})$ and
$\mu_{k}:=\mathbb{E}[a_{t_{k}}\mid\mathcal{F}_{(n)}]\in\Delta(A)$ (27)
denotes the distribution over the action profiles played at the $k$th visit to $s$ conditioned on $\mathcal{F}_{(n)}$. We can invoke Lemma 1 to characterize the distribution $\mu_{k}$ for $k=k_{(n)},\ldots,k_{(n+1)}-1$ within epoch $n$ in comparison to the stationary distribution of the logit dynamics in a fictional (stationary) scenario. To this end, we focus on the evolution of the action profiles across consecutive visits to state $s$ while modeling the state-action trajectories in between the consecutive visits to $s$ as some external (possibly coupled) processes, as illustrated in Figure 1.
Let $Q_{(n)}(\cdot):=Q_{t_{(n)}}(\cdot)$ and $c_{(n)}(\cdot):=c_{t_{(n)}}(\cdot)$. [^5] We drop the agent index $i$ in $Q_{t}^{i}$ since the agents have common estimates. For the fictional scenario, we consider the logit dynamics followed in the repeated play of $\mathcal{G}_{(n)}=\langle(A^{i})_{i\in[N]},Q_{(n)}(s,\cdot)\rangle$ at every visit to state $s$ in the main scenario. Correspondingly, $\mu(Q_{(n)}(s,\cdot))\in\Delta(A)$ is the unique stationary distribution of the action profiles for the logit dynamics followed in $\mathcal{G}_{(n)}$. On the other hand, by (9) and Assumption 1-$(i)$, we can bound the changes on the Q-function estimates within epoch $n$ in the main scenario by
$|Q_{t}(s,a)-Q_{(n)}(s,a)|\leq HK_{(n)}\beta_{(n)}\quad\forall a\in A,\;t\in[t_{(n)},t_{(n+1)}),$ (28)
where we define $\beta_{(n)}:=\max_{a}\{\beta_{c_{(n)}(s,a)}\}$ and $H:=2/(1-\gamma)$ is a uniform upper bound on $|r_{t}-\gamma v_{t}(s_{t+1})-Q_{t}(s,a)|$ for all $(s,a)$ and $t$.
The following lemma characterizes $\mu_{k}$ in terms of $\mu(Q_{(n)}(s,\cdot))$ by coupling the evolution of the dynamics in the main and fictional scenarios based on (28).
Lemma 2.
For sufficiently large $n$, we have
$\|\mu_{k}-\mu(Q_{(n)}(s,\cdot))\|_{1}\leq C\rho^{k-k_{(n)}}+DK_{(n)}\beta_{(n)}\quad\forall k\in[k_{(n)},k_{(n+1)})$ (29)
for some constants $C,D>0$ and $\rho\in(0,1)$.
The proof (provided in Appendix B) follows from Proposition 2 and Lemma 1. We highlight the resemblance of (29) to (24).
By (7), the empirical average $\pi^{\mathrm{ave}}_{t}(s)$ evolves according to
$\pi^{\mathrm{ave}}_{t+1}(s)=\pi^{\mathrm{ave}}_{t}(s)+\mathds{1}_{\{s_{t}=s\}}\cdot\frac{1}{c_{t}(s)}\cdot(a_{t}-\pi^{\mathrm{ave}}_{t}(s)).$ (30)
Henceforth, we view actions $a\subset\Delta(A)$ as pure strategies, or probability vectors, in which the associated action gets played with probability $1$. By (30), $\pi^{\mathrm{ave}}_{t}(s)$ gets updated only when $s$ gets visited, i.e., $\pi^{\mathrm{ave}}_{t}(s)=\pi^{\mathrm{ave}}_{t_{k}}(s)$ for all $t\in[t_{k},t_{k+1})$. Then, we can write (30) as
$\pi^{\mathrm{ave}}_{t_{k+1}}(s)=\pi^{\mathrm{ave}}_{t_{k}}(s)+\alpha_{k}\cdot(a_{t_{k}}-\pi^{\mathrm{ave}}_{t_{k}}(s)),$ (31)
where $\alpha_{k}:=1/k$.
The following lemma shows that accumulated changes can still form a stochastic approximation recursion for certain (growing) epoch lengths. Its proof is available in Appendix C.
Lemma 3.
Consider the iterative update rule (31) with some step sizes $\{\alpha_{k}\in[0,1]\}$ satisfying the standard conditions that $\alpha_{k}\rightarrow 0$ monotonically, $\sum_{k}\alpha_{k}=\infty$ and $\sum_{k}\alpha_{k}^{2}<\infty$. Then, (31) yields that we can write the accumulated changes in $\pi^{\mathrm{ave}}_{(n)}(s):=\pi^{\mathrm{ave}}_{t_{(n)}}(s)$ across epochs $n=1,2,\ldots$ as
$\pi^{\mathrm{ave}}_{(n+1)}(s)=\pi^{\mathrm{ave}}_{(n)}(s)+\overline{\alpha}_{(n)}\cdot\left(\sum_{k=k_{(n)}}^{k_{(n+1)}-1}\frac{\overline{\alpha}_{k}}{\overline{\alpha}_{(n)}}\cdot a_{t_{k}}-\pi^{\mathrm{ave}}_{(n)}(s)\right),$ (32)
where the auxiliary step sizes are defined by
$\overline{\alpha}_{(n)}:=\sum_{k=k_{(n)}}^{k_{(n+1)}-1}\overline{\alpha}_{k}\quad\mbox{and}\quad\overline{\alpha}_{k}:=\alpha_{k}\prod_{l=k+1}^{k_{(n+1)}-1}(1-\alpha_{l})\quad\forall k\in[k_{(n)},k_{(n+1)}).$ (33)
Furthermore, there exist epoch lengths $K_{(n)}:=k_{(n+1)}-k_{(n)}$ growing at a sufficiently slow rate that $K_{(n)}\rightarrow\infty$ and $K_{(n)}/k_{(n)}\rightarrow 0$ as $n\rightarrow\infty$ while the step sizes $\{\overline{\alpha}_{(n)}\in[0,1]\}$ can still satisfy the standard conditions that $\overline{\alpha}_{(n)}\rightarrow 0$, $\sum_{n}\overline{\alpha}_{(n)}=\infty$, and $\sum_{n}\overline{\alpha}_{(n)}^{2}<\infty$.
For $\alpha_{k}=1/k$, the accumulated update (32) can be written as
$\pi^{\mathrm{ave}}_{(n+1)}(s)=\pi^{\mathrm{ave}}_{(n)}(s)+\overline{\alpha}_{(n)}\cdot\left(\mu(Q_{(n)}(s,\cdot))+e_{(n+1)}+\omega_{(n+1)}-\pi^{\mathrm{ave}}_{(n)}(s)\right),$ (34)
where $\overline{\alpha}_{(n)}=K_{(n)}/(k_{(n+1)}-1)$ by (33) and the auxiliary error terms are defined by
$e_{(n+1)}:=\frac{1}{K_{(n)}}\sum_{k=k_{(n)}}^{k_{(n+1)}-1}(\mu_{k}-\mu(Q_{(n)}(s,\cdot))),\quad\omega_{(n+1)}:=\frac{1}{K_{(n)}}\sum_{k=k_{(n)}}^{k_{(n+1)}-1}(a_{t_{k}}-\mu_{k}).$ (35)
Both $e_{(n+1)}$ and $\omega_{(n+1)}$ depend on $s$ implicitly for notational simplicity. The bound on $\|\mu_{k}-\mu(Q_{(n)}(s,\cdot))\|$ in Lemma 2 yields that the error $e_{(n+1)}$ is bounded by
$\left\|e_{(n+1)}\right\|_{1}\leq\frac{C}{1-\rho}\cdot\frac{1}{K_{(n)}}+D\cdot K_{(n)}\beta_{(n)}$ (36)
for sufficiently large $n$, due to the triangle inequality. On the other hand, $\omega_{(n+1)}$ forms a square-integrable Martingale difference sequence by the definition (27).
At the same timescale with (34), we can write the accumulated changes for the Q-update (9) as
$Q_{(n+1)}(s,\cdot)=Q_{(n)}(s,\cdot)+\overline{\alpha}_{(n)}\cdot\epsilon_{(n+1)}$ (37)
for some error term $\epsilon_{(n+1)}\in\mathbb{R}^{|A|}$, bounded by
$\|\epsilon_{(n+1)}\|_{\infty}\leq H\beta_{(n)}k_{(n+1)}$ (38)
due to (28).
The following lemma shows that the error terms are asymptotically negligible.
Lemma 4.
We have $\beta_{(n)}k_{(n+1)}\rightarrow 0$, and therefore, $\|e_{(n)}\|\rightarrow 0$ and $\|\epsilon_{(n)}\|\rightarrow 0$ as $n\rightarrow\infty$ almost surely.
Proof.
Recall that agents revising their actions can take any action with some positive probability uniformly bounded away from zero and any agent revises their actions with some fixed positive probability. Since agents can play any action profile at comparable frequencies, the irreducibility assumption on the underlying stochastic game in the statement of Theorem 1 yields that any state $s$ gets visited infinitely often, i.e., $c_{t}(s)\rightarrow\infty$ as $t\rightarrow\infty$ almost surely. Correspondingly, any pair $(s,a)$ gets realized infinitely often, i.e., $c_{t}(s,a)\rightarrow\infty$ as $t\rightarrow\infty$ almost surely. Therefore, if we follow similar lines with the proof of (Sayin et al., 2022a, Lemma 6.1), Assumption 1-$(iii)$ yields that
$\lim_{t\rightarrow\infty}c_{t}(s)\beta_{c_{t}(s,a)}=0\quad\forall(s,a),$ (39)
with probability $1$.
Recall also that $k_{(n)}=c_{(n)}(s)$ and $c_{(n)}(s,a)$ denote, resp., the number of times $s$ and $(s,a)$ get realized in the first stage of epoch $n$. By (39), $k_{(n)}\beta_{c_{(n)}(s,a)}$ decays to zero for all $(s,a)$, and therefore, $k_{(n)}\beta_{(n)}$ decays to zero since there are finitely many pairs and $\beta_{(n)}=\max_{a}\{\beta_{c_{(n)}(s,a)}\}$.
Let $K_{(n)}$ be chosen according to Lemma 3. Then, such a choice of $K_{(n)}$ satisfies $k_{(n)}\geq n\geq K_{(n)}$ as $\{K_{(n)}\}$ grows at a sub-linear rate and $K_{(n)}\geq 1$ for all $n$. Therefore, $K_{(n)}\beta_{(n)}$ and $k_{(n+1)}\beta_{(n)}=K_{(n)}\beta_{(n)}+k_{(n)}\beta_{(n)}$ decay to zero almost surely. By (36) and (38), the error terms also decay to zero almost surely. ∎
Lemma 4 yields that we can rewrite the accumulated changes (34) and (37), resp., on $\pi^{\mathrm{ave}}_{(n)}(s)$ and $Q_{(n)}(s,\cdot)$ across epochs as
$\pi^{\mathrm{ave}}_{(n+1)}(s)=\pi^{\mathrm{ave}}_{(n)}(s)+\overline{\alpha}_{(n)}\cdot(\mu(Q_{(n)}(s,\cdot))-\pi^{\mathrm{ave}}_{(n)}(s)+\omega_{(n+1)}+\mathrm{a.n.e.})$ (40a)
$Q_{(n+1)}(s,\cdot)=Q_{(n)}(s,\cdot)+\overline{\alpha}_{(n)}\cdot\mathrm{a.n.e.},$ (40b)
where $\mathrm{a.n.e.}$ refers to asymptotically negligible error terms. Since step sizes $\overline{\alpha}_{(n)}$ satisfies the stochastic approximation conditions by Lemma 3, the function $\mu(\cdot)$ is locally Lipschitz continuous by Proposition 2, $\omega_{(n+1)}$ forms a square-integrable Martingale difference sequence, and the iterates $(\pi^{\mathrm{ave}}_{(n)}(s),Q_{(n)}(s,\cdot))$ are bounded, we can approximate the discrete-time accumulated updates (40) with the following continuous-time flow
$\dot{\pi}=\mu(Q)-\pi\quad\mbox{and}\quad\dot{Q}=0$ (41)
such that (40) converges to an internally chain-recurrent set of (41) (Benaim, 1999). Furthermore, we can characterize its internally chain-recurrent set by formulating a Lyapunov function (Benaim, 1999). For example, the Lyapunov function $V(\pi,Q)=\|\mu(Q)-\pi\|_{2}^{2}$ yields that
$\lim_{n\rightarrow 0}\|\pi^{\mathrm{ave}}_{(n)}(s)-\mu(Q_{(n)}(s,\cdot))\|_{1}=0$ (42)
almost surely as norms are comparable in finite-dimensional spaces.
We have the tracking result (42) for the first stages $t_{(n)}$ of each epoch $n$. We can also show the tracking result for each $t$ since
$\|\pi^{\mathrm{ave}}_{t}(s)-\mu(Q_{t}(s,\cdot))\|_{1}\leq$ $\;\|\pi^{\mathrm{ave}}_{t}(s)-\pi^{\mathrm{ave}}_{(n)}(s)\|_{1}+\|\mu(Q_{(n)}(s,\cdot))-\mu(Q_{t}(s,\cdot))\|_{1}$
$+\|\pi^{\mathrm{ave}}_{(n)}(s)-\mu(Q_{(n)}(s,\cdot))\|_{1}$
for all $t\in[t_{(n)},t_{(n+1)})$, by the triangle inequality. The first term on the right-hand side is bounded by
$\|\pi_{t}(s)-\pi_{(n)}(s)\|_{1}\leq 2K_{(n)}/{k_{(n)}}\quad\forall t\in[t_{(n)},t_{(n+1)})$ (43)
by (30) and decays to zero by Lemma 3. The second one is bounded by
$\|\mu(Q_{t}(s,\cdot))-\mu(Q_{(n)}(s,\cdot))\|_{1}\leq LHK_{n}\beta_{(n)}\quad\forall t\in[t_{(n)},t_{(n+1)})$ (44)
for some constant $L$, by Proposition 2 and (28), and decays to zero by Lemma 4. The last one decays to zero by (42). This completes the proof of Proposition 4 since the fixed state $s$ is arbitrary. ∎
5.3 Convergence of the Q-function Estimates
The convergence of the Q-function estimates to $Q_{*}$, as described in (12), follows from showing that the value function estimates track the joint maximizer of the Q-function estimates approximately, as shown in the following corollary to Proposition 4.
Corollary 3.
We have $\limsup_{t\rightarrow\infty}|v_{t}(s)-\max_{a}\{Q_{t}(s,a)\}|\leq e_{Q}\cdot(1-\gamma)/\gamma$ almost surely, where $e_{Q}\geq 0$ is as described in Table 2.
Proof.
For AL-Q, we have $\pi_{t}\equiv\pi^{\mathrm{ave}}_{t}$ and $\|\pi_{t}(s)-\mu^{\mathrm{coor}}(Q_{t}(s,\cdot))\|_{1}=\|\pi_{t}(s)-\underline{\mathrm{br}}(Q_{t}(s,\cdot))\|_{1}\rightarrow 0$ as $t\rightarrow\infty$ by Propositions 3 and 4. This implies that
$\limsup_{t\rightarrow\infty}|v_{t}(s)-\max_{a}\{Q_{t}(s,a)\}|\leq\tau\log|A|$ (45)
since $|\max_{a}\{Q_{t}(s,a)\}-\mathbb{E}_{a\sim\underline{\mathrm{br}}(Q_{t}(s,\cdot))}[Q_{t}(s,a)]\,|\leq\tau\log|A|$.
For AIL-Q, we have $\pi_{t}\equiv\pi^{\mathrm{ave}}_{t}$ and $\|\pi_{t}(s)-\mu^{\mathrm{ind}}(Q_{t}(s,\cdot))\|_{1}\rightarrow 0$ as $t\rightarrow\infty$ by Proposition 4. By the triangle inequality, we have
$|v_{t}(s)-\max_{a}\{Q_{t}(s,a)\}|\leq$ $\;|\mathbb{E}_{a\sim\pi_{t}(s)}[Q_{t}(s,a)]-\mathbb{E}_{\mu^{\mathrm{ind}}(s)}[Q_{t}(s,a)]|$
$|\mathbb{E}_{\mu^{\mathrm{ind}}(s)}[Q_{t}(s,a)]-\mathbb{E}_{\mu^{\mathrm{coor}(s)}}[Q_{t}(s,a)]|$
$|\mathbb{E}_{\mu^{\mathrm{coor}(s)}}[Q_{t}(s,a)]-\max_{a}\{Q_{t}(s,a)\}|.$ (46)
The first term on the right-hand side decays to zero while the last is bounded by $\tau\log|A|$. If we invoke Proposition 3 for the second term, we obtain
$\limsup_{t\rightarrow\infty}|v_{t}(s)-\max_{a}\{Q_{t}(s,a)\}|\leq\tau\log|A|+\frac{\Lambda(\delta)}{1-\gamma}$ (47)
since the Q-function estimates are bounded by $1/(1-\gamma)$.
For EIL-Q, we have $\pi_{t}\equiv\pi^{\mathrm{free}}_{t}$ and
$\|\pi^{\mathrm{ave}}_{t}(s)-\underline{\mathrm{br}}(Q_{t}(s,\cdot))\|\leq$ $\;\|\pi^{\mathrm{ave}}_{t}(s)-\mu^{\mathrm{ind}}(Q_{t}(s,\cdot))\|+\|\mu^{\mathrm{ind}}(Q_{t}(s,\cdot))-\mu^{\mathrm{coor}}(Q_{t}(s,\cdot))\|$ (48)
for any norm by the triangle inequality. Then, Propositions 4 and 3 yield
$\limsup_{t\rightarrow\infty}\|\pi^{\mathrm{ave}}_{t}(s)-\underline{\mathrm{br}}(Q_{t}(s,\cdot))\|_{\infty}\leq\Lambda(\delta).$ (49)
as $\|\cdot\|_{\infty}\leq\|\cdot\|_{1}$. Let $a^{1}\in\operatorname*{argmax}_{a}\{\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a)\}$ and $a^{2}\in\operatorname*{argmax}_{a}\{\pi^{\mathrm{ave}}_{t}(s)(a)\}$. Then, we can bound the difference of the probabilities of these action profiles in $\underline{\mathrm{br}}(Q_{t}(s,\cdot))$ by
00 0 $\leq\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a^{1})-\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a^{2})$
$\leq\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a^{1})-\pi^{\mathrm{ave}}_{t}(s)(a^{1})+\pi^{\mathrm{ave}}_{t}(s)(a^{2})-\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a^{2})$
$\leq 2\Lambda(\delta)+\mathrm{a.n.e.},$ (50)
where the first and second inequalities follow since $a^{1}$ and $a^{2}$ are, resp., the maximizers of the distributions $\underline{\mathrm{br}}(Q_{t}(s,\cdot))$ and $\pi^{\mathrm{ave}}_{t}(s)$, and the last inequality follows from (49).
The boundedness of the Q-iterates by $1/(1-\gamma)$ and the smoothed best response (6) guarantee that
$\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a)\geq\underline{\epsilon}:=\frac{1}{|A|}\exp\left(-\frac{1}{\tau(1-\gamma)}\right)\quad\forall(s,a).$ (51)
Therefore, we can bound the difference of these action profiles in $Q_{t}(s,\cdot)$ by
$Q_{t}(s,a^{1})-Q_{t}(s,a^{2})$ $=\tau\cdot\log\left(\frac{\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a^{1})}{\underline{\mathrm{br}}(Q_{t}(s,\cdot))(a^{2})}\right)$ (52)
$\leq\tau\cdot\log\left(1+2\Lambda(\delta)/\underline{\epsilon}+\mathrm{a.n.e.}\right),$ (53)
where the first line follows from the definition (6) and the second one follows from (5.3) and (51). Since $\log(\cdot)$ is a monotonically increasing continuous function, (53) yields that
$\limsup_{t\rightarrow\infty}\{Q_{t}(s,a^{1})-Q_{t}(s,a^{2})\}\leq\Xi(\delta)\coloneqq\tau\cdot\log\left(1+2\Lambda(\delta)/\underline{\epsilon}\right).$ (54)
The inequality (54) holds for any $a^{2}\in A_{t}(s)$ since $A_{t}(s)$, as described in (26), is the set of maximizers of $\pi^{\mathrm{ave}}_{t}(s)$. Moreover, by the definition (6), $Q_{t}(s,a^{1})=\max_{a}\{Q_{t}(s,a)\}$. Since any convex combination of $Q_{t}(s,a^{2})$ over $a^{2}\in A_{t}$, including $\pi^{\mathrm{free}}_{t}(s)\in\Delta(A)$, also satisfies the inequality (54), we obtain
$\limsup_{t\rightarrow\infty}|v_{t}(s)-\max_{a}\{Q_{t}(s,a)\}|\leq\Xi(\delta)$ (55)
almost surely, as $v_{t}(s)=\mathbb{E}_{a\sim\pi^{\mathrm{free}}_{t}(s)}[Q_{t}(s,a)]$.
For EL-Q, the right-hand side of (49) is 00 by Proposition 4. Hence, (55) yields that $|v_{t}(s)-\max_{a}\{Q_{t}(s,a)\}-v_{t}(s)|$ decay to zero almost surely. ∎
We can write the Q-update (9) as
$Q_{t+1}(s,a)=$ $\;Q_{t}(s,a)+\mathds{1}_{\{(s,a)=(s_{t},a_{t})\}}\beta_{c_{t}(s,a)}$
×$\times\Big{(}r(s,a)+\gamma\sum_{s^{\prime}}p(s^{\prime}\mid s,a)\max_{\tilde{a}}\{Q_{t}(s^{\prime},\tilde{a})\}-Q_{t}(s,a)+\bar{\epsilon}_{t}(s,a)+\bar{\omega}_{t+1}(s,a)\Big{)},$ (56)
where the error term and the stochastic approximation noise are, resp., defined by
$\bar{\epsilon}_{t}(s,a)=\gamma\sum_{s^{\prime}}p(s^{\prime}\mid s,a)\big{(}v_{t}(s^{\prime})-\max_{\tilde{a}}\{Q_{t}(s^{\prime},\tilde{a})\}\big{)}$ (57a)
$\bar{\omega}_{t+1}(s,a)=\gamma\left(v_{t}(s_{t+1})-\sum_{s^{\prime}}p(s^{\prime}\mid s,a)v_{t}(s^{\prime})\right).$ (57b)
Corollary 3 and (57a) yield that the error term $\bar{\epsilon}_{t}$ satisfies
$\limsup_{t\rightarrow\infty}|\bar{\epsilon}_{t}(s,a)|\leq(1-\gamma)e_{Q}\quad\forall(s,a)$ (58)
almost surely. On the other hand, since $s_{t+1}\sim p(\cdot\mid s_{t},a_{t})$, the stochastic approximation noise $\bar{\omega}_{t+1}$ forms a square-integrable Martingale difference sequence. Due to the exploration inherent to the logit-Q dynamics, the irreducibility assumption in the statement of Theorem 1 yields that each $(s,a)$ pair gets realized infinitely often. Based on the contraction property of the Bellman optimality operator, Theorem 5.1 in the extended version of (Sayin et al., 2022a) yields (11). [^6] The extended version of (Sayin et al., 2022a) is available at https://arxiv.org/abs/2010.04223.
5.4 Convergence of the Estimated Plays to Equilibrium
Recall that $v_{\pi}:S\rightarrow\mathbb{R}$ and $Q_{\pi}:S\times A\rightarrow\mathbb{R}$, as described in (3), resp., denote the value and Q-function of the Markov stationary policy $\pi:S\rightarrow\Delta(A)$. For example, we have $v_{\pi}(s)=\mathrm{E}_{a\sim\pi(s)}[Q_{\pi}(s,a)]$. Based on $Q_{*}$, as described in (12), we define $v_{*}(s)=\max_{a}\{Q_{*}(s,a)\}$ for all $s$. Then, given the strategy $\pi_{t}$, we have
00 0 $\leq v_{*}(s)-v_{\pi_{t}}(s)=v_{*}(s)-\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)]+\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)-Q_{\pi_{t}}(s,a)].$ (59)
Since $(v_{*},Q_{*})$ and $(v_{\pi_{t}},Q_{\pi_{t}})$ pairs satisfy the one-step Bellman equation, we also have
$\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)-Q_{\pi_{t}}(s,a)]$ $=\gamma\mathrm{E}_{a\sim\pi_{t}(s)}\left[\sum_{s^{\prime}\in S}p(s^{\prime}\mid s,a)(v_{*}(s^{\prime})-v_{\pi_{t}}(s^{\prime}))\right]\leq\gamma\|v_{*}-v_{\pi_{t}}\|_{\infty}$ (60)
for both $\pi_{t}\equiv\pi^{\mathrm{ave}}_{t}$ and $\pi_{t}\equiv\pi^{\mathrm{free}}_{t}$. On the other hand, we can bound the first two terms $v_{*}(s)-\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)]$ on the right-hand side of (59) by
$|v_{*}(s)-$ $\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)]|\leq|\max_{a}\{Q_{*}(s,a)\}-\max_{a}\{Q_{t}(s,a)\}|$
$+|\max_{a}\{Q_{t}(s,a)\}-v_{t}(s)|+|\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)]-\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{t}(s,a)]|$ (61)
Then, we can bound the first and third terms on the right-hand side by (11) and bound the second term by Corollary 3, and obtain
$\limsup_{t\rightarrow\infty}|v_{*}(s)-\mathrm{E}_{a\sim\pi_{t}(s)}[Q_{*}(s,a)]|\leq\frac{1+\gamma}{\gamma}\cdot e_{Q}$ (62)
due to the sub-additivity of the limit superior. Since the upper bound in (62) does not depend on $(s,a)$, (59), (60), and (62) yields that
$\limsup_{t\rightarrow\infty}\|v_{*}-v_{\pi_{t}}\|_{\infty}\leq\frac{1+\gamma}{\gamma(1-\gamma)}e_{Q}.$ (63)
This completes the proof of Theorem 1.
(a) Q-function distance to $Q_{*}$
(b) Value function distance to $v_{*}$
(c) Performance of the average play $\pi_{t}$
(d) Q-function distances of agents
Figure 2: Performance of the AL-Q, EL-Q, AIL-Q, and EIL-Q after $10$ independent runs. From top to bottom, the solid red, green, blue, and magenta curves represent the AL-Q, EL-Q, AIL-Q, and EIL-Q dynamics, respectively. The shaded areas around the solid curves represent the standard deviation across the independent runs. In Figures 2(a), 2(b), and 2(c), all four dynamics converge to the associated efficient equilibrium values, corroborating Theorem 1. Furthermore, in Figure 2(d), the impact of different initializations on the long-run behavior of the dynamics diminishes in time, showing their robustness to arbitrary initializations.
6 Illustrative Examples
In this section, we examine the long-run behavior of Algorithm 1 numerically. Consider a stochastic team $\langle S,\{A^{i}\}_{i\in[N]},r,p,\gamma\rangle$, where there are three states, i.e., $|S|=3$, and three agents, i.e., $N=3$. Each agent can take two actions in each state, i.e., $|A^{i}|=2$ for each $i$ and $|A|=8$. The stage-payoffs are randomly selected such that Q-function estimates remain bounded within the $[0,1]$ interval. We set $\gamma=0.6$, and $\tau=0.05$. Furthermore, we set $\delta=0.2$ for the independent-revision schemes. For all three agents, the initial Q-function values are assigned differently, e.g., $Q_{0}^{i}(s,a)=1+0.1\cdot i$ for all $(s,a)$, so that we can examine the impact of different initializations. We run the experiments for $10^{6}$ time steps over $10$ independent trials. Figure 2 depicts both the mean error and standard deviation around the mean across these trials for the AL-Q, EL-Q, AIL-Q, and EIL-Q algorithms. We use logarithmic scale for the horizontal axis, i.e., for the time, to illustrate the transient and steady-state behavior of the dynamics together more explicitly.
In Figures 2(a) and 2(b), we examine the convergence of the Q-function estimates $Q_{t}$ and the value function estimates $v_{t}$ to the optimal values $Q_{*}$ and $v_{*}$, respectively. We observe that the estimates in all four algorithms converge to the ones associated with the efficient equilibrium, implying that the logit-Q dynamics can learn efficient equilibrium in stochastic teams. However, the independent-revision schemes are relatively more noisy in early stages than the coordinated ones.
Figure 2(c) illustrates the performance of the estimated play for the AL-Q, EL-Q, AIL-Q, and EIL-Q algorithms. More specifically, the estimated plays $\pi^{\mathrm{ave}}_{t}$’s in the averaging approaches from the AL-Q, and AIL-Q dynamics converge to near efficient equilibrium due to the inherent exploration caused by the smoothed best responses while the estimated plays $\pi^{\mathrm{free}}_{t}$’s in the exploration-free approaches from the EL-Q, and EIL-Q dynamics converge exactly to the efficient equilibrium.
Lastly, we empirically check whether uncoordinated initializations of agents’ $Q$-function estimates cause non-diminishing differences in their estimates in Figure 2(d). We observe that when agents initialize their $Q$-function estimates arbitrarily, the differences in the $Q$-function estimates across agents diminish in time.
7 Concluding Remarks
This paper introduced a new family of logit-Q dynamics for efficient learning in unknown stochastic teams by combining log-linear learning with Q-learning within the stage-game framework. We proposed four new dynamics: Averaged Logit-Q (AL-Q), Exploration-free Logit-Q (EL-Q), Averaged Independent Logit-Q (AIL-Q), and Exploration-free Independent Logit-Q (EIL-Q). These dynamics differ in terms of how agents revise their actions and model their joint play, offering flexibility for different learning scenarios. We showed that these dynamics converge to a (near) efficient equilibrium in stochastic teams, with quantifiable approximation errors. We also demonstrated the rationality of the dynamics against agents following pure stationary strategies and stochastic games with a single controller where the stage-payoffs specific to each state induces a potential game. We conducted various simulations to illustrate the convergence numerically. Future research could explore scaling these dynamics using function approximation or networked interconnections.
Acknowledgment
This work was supported by The Scientific and Technological Research Council of Türkiye (TUBITAK) BIDEB 2232-B International Fellowship for Early Stage Researchers under Grant Number 121C124.
Appendix A Proof of Lemma 1
The difference between the distributions $\mu_{k}^{w}$ and $\widehat{\mu}_{k}$ is bounded by $\|\mu_{k}^{w}-\widehat{\mu}_{k}\|_{1}\leq 2\Pr(w_{k}\neq\widehat{w}_{k})$ due to the coupling lemma, e.g., see (Levin and Peres, 2017, Proposition 4.2) and (Levin and Peres, 2017, Proposition 4.7). To bound $\Pr(w_{k}\neq\widehat{w}_{k})$, we construct a coupling over $\left\{\widehat{w}_{k}\right\}$, and $\left\{(w_{k},z_{k})\right\}$ based on the following two cases:
Case $(i)$. The realized states of the processes do not match, i.e., $w_{k}\neq\widehat{w}_{k}$. Then, the processes $\{(w_{k},z_{k})\}_{k\geq 0}$ and $\{\widehat{w}_{k}\}_{k\geq 0}$, resp., transit to $(w_{k+1},z_{k+1})$ and $\widehat{w}_{k+1}$ independently according to their transition kernels $\mathbb{P}$ and $\widehat{\mathbb{P}}$.
Case $(ii)$. The realized states of the processes match, i.e., $w_{k}=\widehat{w}_{k}$. Then, the processes $\{(w_{k},z_{k})\}_{k\geq 0}$ and $\{\widehat{w}_{k}\}_{k\geq 0}$, resp., transit to $(w_{k+1},z_{k+1})$ and $\widehat{w}_{k+1}$ still according to their transition kernels $\mathbb{P}$ and $\widehat{\mathbb{P}}$, but in a coupled way such that we have
$\Pr\left(w_{k+1}\neq\widehat{w}_{k+1}\mid\widehat{w}_{k}=w_{k},\{w_{l},z_{l}\}_{l=0}^{k}\right)=\left\|\mathbb{P}^{w}(\cdot\mid\{w_{l},z_{l}\}_{l\leq k})-\widehat{\mathbb{P}}(\cdot\mid w_{k})\right\|_{\mathrm{TV}}$ (64)
for any $\{(w_{l},z_{l})\in W\times Z\}_{l\leq k}$. Such a coupling is possible based on (Levin and Peres, 2017, Proposition 4.7).
Condition $(ii)$ and Case $(ii)$ in the coupling constructed yield that
$\Pr\left(w_{k+1}\neq\widehat{w}_{k+1}\mid\widehat{w}_{k}=w_{k},\{w_{l},z_{l}\}_{l=0}^{k}\right)\leq 1-\lambda\quad\forall\{(w_{l},z_{l})\in W\times Z\}_{l\leq k}$ (65)
for any $t$. Moreover, the following is true
$\Pr\Big{(}w_{k+1}\neq\widehat{w}_{k+1}\mid\widehat{w}_{k}=w_{k},\{w_{l}\}_{l=0}^{k}\Big{)}$ =$=\sum_{\{z_{l}\}_{l\leq k}\in Z^{k}}\Big{(}\Pr\left(w_{k+1}\neq\widehat{w}_{k+1}\mid\widehat{w}_{k}=w_{k},\{w_{l},z_{l}\}_{l=0}^{k}\right)$
×$\hskip 85.35826pt\times\Pr(\{z_{l}\}_{l\leq k}\mid\widehat{w}_{k}=w_{k},\{w_{l}\}_{l=0}^{k})\Big{)}$
$\leq\sum_{\{z_{l}\}_{l\leq k}\in Z^{k}}(1-\lambda)\Pr(\{z_{l}\}_{l\leq k}\mid\widehat{w}_{k}=w_{k},\{w_{l}\}_{l=0}^{k})$
$=1-\lambda,\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}$ (66)
for all $k$, and$\{w_{l}\}_{l=0}^{k}$.
On the other hand, by Condition $(i)$, for any $\widehat{w}_{k}\neq w_{k}$, there exists $\kappa$-length paths $\{\widehat{w}_{k+l}=\widehat{h}^{l}\}_{l=0}^{\kappa}$ and $\{w_{k+l}=h^{l}\}_{l=0}^{\kappa}$ where $\widehat{h}^{l}\neq h^{l}$ for all $l=0,\ldots,\kappa-1$ and $\widehat{h}^{\kappa}=h^{\kappa}$. Therefore, for the states that are $\kappa$ stages away given that the processes do not visit the same state initially, we have
$\Pr(w_{k+\kappa}=\widehat{w}_{k+\kappa}\mid w_{k}\neq\widehat{w}_{k})$ $\geq\Pr\left(\bigwedge_{k=1}^{\kappa}(\widehat{w}_{k+l},w_{k+l})=(\widehat{h}^{l},h^{l})\mid\widehat{w}_{k}\neq w_{k}\right)\geq\varepsilon\widehat{\varepsilon},$ (67)
where the second inequality follows from (22) and Case $(i)$ of the coupling constructed as the processes evolve independently when they do not match and they do not match along these paths until $k+\kappa$.
Partition the time into $\kappa$-length intervals. If $w_{(m-1)\kappa}\neq\widehat{w}_{(m-1)\kappa}$, (67) yields that
$\Pr(w_{m\kappa}=\widehat{w}_{m\kappa}\mid w_{(m-1)\kappa}\neq\widehat{w}_{(m-1)\kappa})\geq\varepsilon\widehat{\varepsilon}.$ (68)
On the other hand, if $w_{(m-1)\kappa}=\widehat{w}_{(m-1)\kappa}$, we have
$\Pr(w_{m\kappa}=\widehat{w}_{m\kappa}\mid w_{(m-1)\kappa}=\widehat{w}_{(m-1)\kappa})$ $\geq\prod_{t=(m-1)\kappa+1}^{m\kappa}\Pr\left(w_{t}=\widehat{w}_{t}\mid\bigwedge_{\tau=(m-1)\kappa}^{t-1}w_{\tau}=\widehat{w}_{\tau}\right)\geq\lambda^{\kappa}$ (69)
for any $m=1,2,\ldots$, where the last inequality follows from (A).
By the bounds on the conditional probabilities (68) and (69), we can bound the unconditional probability by
$\Pr(w_{m\kappa}\neq\widehat{w}_{m\kappa})=$ $\;\Pr(w_{m\kappa}\neq\widehat{w}_{m\kappa}\mid w_{(m-1)\kappa}\neq\widehat{w}_{(m-1)\kappa})\times\Pr(w_{(m-1)\kappa}\neq\widehat{w}_{(m-1)\kappa})$
$+\Pr(w_{m\kappa}\neq\widehat{w}_{m\kappa}\mid w_{(m-1)\kappa}=\widehat{w}_{(m-1)\kappa})\times\Pr(w_{(m-1)\kappa}=\widehat{w}_{(m-1)\kappa}).$
$\leq$ $\;(1-\varepsilon\widehat{\varepsilon})\times\Pr(w_{(m-1)\kappa}\neq\widehat{w}_{(m-1)\kappa})+1-\lambda^{\kappa}.$ (70)
Hence, by induction, we obtain
$\Pr(w_{m\kappa}\neq\widehat{w}_{m\kappa})$ $\leq(1-\varepsilon\widehat{\varepsilon})^{m}+\frac{1-\lambda^{\kappa}}{\varepsilon\widehat{\varepsilon}}\quad\forall m=0,1,2,\ldots.$ (71)
Lastly, for any $k\in[m\kappa,(m+1)\kappa)$ and $m=0,1,\ldots$, the mismatch probability is bounded from above by
$\Pr(w_{k}\neq\widehat{w}_{k})=$ $\;\Pr(w_{k}\neq\widehat{w}_{k}\mid w_{m\kappa}\neq\widehat{w}_{m\kappa})\times\Pr(w_{m\kappa}\neq\widehat{w}_{m\kappa})$
$+(1-\Pr(w_{k}=\widehat{w}_{k}\mid w_{m\kappa}=\widehat{w}_{m\kappa}))\times\Pr(w_{m\kappa}=\widehat{w}_{m\kappa})$
$\leq$ $\;\Pr(w_{m\kappa}\neq\widehat{w}_{m\kappa})+1-\lambda^{\kappa}$ (72)
since $\Pr(w_{k}=\widehat{w}_{k}\mid w_{m\kappa}=\widehat{w}_{m\kappa})\geq\lambda^{k-m\kappa}\geq\lambda^{\kappa}$ as in (69). Combining (71) and (72), we obtain
$\Pr\left(w_{k}\neq\widehat{w}_{k}\right)\leq(1-\varepsilon\widehat{\varepsilon})^{m}+(1-\lambda^{\kappa})\frac{1+\varepsilon\widehat{\varepsilon}}{\varepsilon\widehat{\varepsilon}}\quad\forall k\in[m\kappa,(m+1)\kappa).$ (73)
for any $m=0,1,\ldots$. Since $1-\varepsilon\widehat{\varepsilon}\in(0,1)$, we have $(1-\varepsilon\widehat{\varepsilon})^{m}\leq(1-\varepsilon\widehat{\varepsilon})^{k/\kappa-1}$ for all $k\in[m\kappa,(m+1)\kappa)$. This completes the proof.
Appendix B Proof of Lemma 2
Recall that we limit our focus to the timesteps in which the state $s$ is visited. Let us denote the time index for the $k$-th visit to state $s$ as $t_{k}$, and the action profile played at the $k$-th visit to state $s$ as $w_{k}=(s,a_{t_{k}})$. Similarly, let us denote the action profile played at the $k$-th visit to state $s$ in the fictional scenario as $\widehat{w}_{k}=(s,\widehat{a}_{t_{k}})$. Correspondingly, let $\mu_{k}^{w}\in\Delta(W)$ and $\widehat{\mu}_{k}\in\Delta(W)$ denote, resp., the distribution of $w_{k}$ and $\widehat{w}_{k}$ at stage $t_{k}$ conditioned on $\mathcal{F}_{(n)}$.
The probability of transition from the action $\widehat{w}$ to $\widehat{w}^{\prime}$ depends on $\widehat{w}$ and $Q_{(n)}$. Since $Q_{(n)}$ is $\mathcal{F}_{(n)}$-measurable, $\{\widehat{w}_{k}\}_{k}$ form a time homogeneous Markov chain conditioned on $\mathcal{F}_{(n)}$. Let $\widehat{\mathbb{P}}$ denote the transition matrix of this time homogeneous Markov chain. On the other hand, different from $\{\widehat{w}_{k}\}_{k}$, the process $\{w_{k}\}_{k}$ does not form a time homogeneous Markov chain since the probability of transition to the next state $w_{k+1}$ is determined not only by the last action profile $w_{k}$, but also by the current $Q$-function estimate $Q_{t_{k+1}}(s,\cdot)$ which is not $\mathcal{F}_{(n)}$-measurable. Recall that the trajectory of state-action pairs realized in between the $k$th and $(k+1)$st visit to state $s$ is denoted as $z_{k}\in\mathbb{Z}$, where $\mathbb{Z}$ represents the countable set of all possible state-action pair trajectories that can be realized in-between two consecutive visits to state $s$. Accordingly, the estimate $Q_{t_{k+1}}(s,a)$ is $\sigma(\{w_{l},z_{l}\}_{l\leq k})$-measurable for all $(s,a)$ due to the Q-update (9). Therefore, the estimate $Q_{(n)}$ at the beginning of the $n$th epoch which is $\mathcal{F}_{(n)}$-measurable, the history of state and action pairs $\{w_{l},z_{l}\}_{l<k}$ from the $(k_{(n)}-1)$st visitation to the $k>(k_{(n)}-1)$st visitation to state $s$ determine the transition probability from $w_{k-1}=(s,a_{t_{k-1}})$ to $w_{k}=(s,a_{t_{k}})$. Let ($\mathbb{P}(\cdot\mid\{w_{l},z_{l}\}_{l\leq k})$ denote the transition probabilities governing the evolution of $\{(w_{k},z_{k})\}$. This implies that the discrete-time processes $\{\widehat{w}_{k}\}$ and $\{(w_{k},z_{k})\}$ are in the framework of Lemma 1. Correspondingly, in the following, we prove the claims that the conditions listed in Lemma 1 hold so that we can invoke the lemma to address the distance between $\mu_{k}^{w}$ and $\widehat{\mu}_{k}$.
Claim 1.
The processes $\{\widehat{w}_{k}\}$ and $\{(w_{k},z_{k})\}$ satisfy Condition $(i)$ in Lemma 1 for $\kappa=|A|$, $\varepsilon=(\epsilon\delta(1-\delta)^{n-1})^{2\kappa}$, and any $w^{\prime}\in W$.
Proof.
Due to the update rule of log-linear learning $\widehat{w}=(s,\widehat{a})\in W$ can transit to $\widehat{w}=(s,\widehat{a}^{\prime})\in W$ only if $\widehat{a}$, and $\widehat{a}^{\prime}$ are
  • arbitrarily different at most for one agent’s action in coordinated scheme,
  • arbitrarily different for every agent’s actions in independent scheme.
This yields that the Markov chain $\{\widehat{w}_{k}\}_{k\geq k_{(n)}}$ is irreducible in both cases. For example, let $n=2$, $A^{1}=\{a,b\}$, and $A^{2}=\{\bar{a},\bar{b}\}$; and consider the action profiles $\widehat{w}=(s,a,\bar{a})$ and $\widehat{w}^{\prime}=(s,b,\bar{b})$ at state $s$. The Markov chain can follow the path $(s,a,\bar{a})$, $(s,a,\bar{b})$, and $(s,b,\bar{b})$, from $\widehat{w}$ to $\widehat{w}^{\prime}$ with some positive probability. The Markov chain also has self-loops for every extended state.
Since the Markov chain $\{\widehat{w}_{k}\}$ is irreducible and has self loops at every state, there exists a $\kappa$-length path $\widehat{w}=\widehat{h}_{0},\ldots,\widehat{h}_{\kappa}=\widehat{w}^{\prime}$ for any $(\widehat{w},\widehat{w}^{\prime})\in W\times W$ in the fictional scenario, where $\kappa=|A|$. Even though $\{(w_{k},z_{k})\}$ do not form a homogeneous Markov chain, $\widehat{\mathbb{P}}$ and $\mathbb{P}$ have the same pattern. In other words, given any trajectory $\{w_{l},z_{l}\}_{l<k}$, we have $\mathbb{P}^{w}(w^{\prime}\mid,\{w_{l},z_{l}\}_{l\leq k})>0$ if and only if ^$\widehat{\mathbb{P}}(w^{\prime}\mid w_{k})>0$ since agents updating their actions can take any action with some positive probability irrespective of the Q-function estimate due to the logit response (6) and the main and fictional scenarios differ only in terms of the Q-function estimates used. Therefore, there also exists a $\kappa$-length path $w=h^{0},\ldots,h^{\kappa}=w^{\prime}$ for any $(w,w^{\prime})\in W\times W$ in the main scenario.
In coordinated scheme, the probabilities of every feasible transitions of the extended states in both main and fictional scenarios are uniformly bounded from below by $\epsilon/n>0$ irrespective of the trajectory followed. In the lower bound, $1/n$ corresponds to the probability that one specific agent updates his/her action, and $\epsilon>0$ is the lower bound on the probability of actions played according to the logit choice as described in (51). On the other hand, the feasible transitions for coordinated scheme are also feasible for independent scheme. In independent scheme, irrespective of the trajectory, such transitions have probabilities that are uniformly bounded from below by $\epsilon\delta(1-\delta)^{n-1}>0$. Here, $\delta(1-\delta)^{n-1}$ corresponds to the probability that only one specific agent updates his/her action due to the independent update scheme. Note that $\delta(1-\delta)^{n-1}<1/n$. Therefore, the $\kappa$-length paths $\{\widehat{w}_{k}=\widehat{h}_{k}\}_{k=0}^{\kappa}$ and $\{w_{k}=h_{k}\}_{k=0}^{\kappa}$ (that are feasible for both coordinated and independent schemes) have probabilities greater than $(\epsilon\delta(1-\delta)^{n-1})^{\kappa}$ in both main and fictional scenarios, i.e.,
$\Pr\left(\bigwedge_{l=1}^{\kappa}\widehat{w}_{k+l}=\widehat{h}_{l}\mid\widehat{a}_{t_{k}}=\widehat{h}_{0}\right)\geq(\epsilon\delta(1-\delta)^{n-1})^{\kappa},$ (74a)
$\Pr\left(\bigwedge_{l=1}^{\kappa}w_{k+l}=h_{l}\mid(w_{k},z_{k})=(h_{0},z)\right)\geq(\epsilon\delta(1-\delta)^{n-1})^{\kappa}\quad\forall z\in\mathbb{Z}.$ (74b)
Given some $w^{\prime}\in W$, consider any pair $(\widehat{w},w)\in W\times W$ such that $\widehat{w}\neq w$. Based on the patterns of the transition matrices $\widehat{\mathbb{P}}$ and $\mathbb{P}$, let ^$\widehat{\ell}\leq\kappa$ and $\ell\leq\kappa$ be the lengths of the shortest paths where action profiles differ at most for one agent’s action (so that it is feasible for both coordinated and independent schemes). We denote these paths, resp., by $\widehat{w}=\widehat{h}_{0},\ldots,\widehat{h}_{\kappa}=w^{\prime}$ and $w=h_{0},\ldots,h_{\kappa}=w^{\prime}$. Without loss of generality, we can assume that if these shortest paths have a common state, they are identical after the first common state since they are the shortest paths to a common destination $a^{\prime}$ and the underlying transitions have the same pattern. If $\ell<\widehat{\ell}$ (or $\widehat{\ell}<\ell$), let $\{w_{k}\}$ (or $\{\widehat{w}_{k}\}$) follow the $\ell$-length (or $\widehat{\ell}$-length) shortest path and then have self-loops at $a^{\prime}$ for $\kappa-\ell$ (or $\kappa-\widehat{\ell}$) times while $\{\widehat{w}_{k}\}$ (or $\{w_{k}\}$) has self-loops at $\widehat{w}$ (or $w$) for $\kappa-\widehat{\ell}$ (or $\kappa-\ell$) times and then follow the $\widehat{\ell}$-length (or $\ell$-length) shortest path. If $\ell=\widehat{\ell}$, then ,^$\ell,\widehat{\ell}<\kappa$ since =^$\ell=\widehat{\ell}=\kappa$ implies that both path traverse every state, which contradicts with the fact that they are the shortest ones and $\widehat{w}\neq w$. Let $\{w_{k}\}$ follow the $\ell$-length shortest path and then have self loops at $w^{\prime}$ for $\kappa-\ell$ times while $\{\widehat{a}_{t_{k}}\}$ has self-loops at $\widehat{a}$ for $\kappa-\ell$ times and then follow the $\ell$-length shortest path. In either case, the processes do not visit the same state concurrently for $k=0,\ldots,\kappa-1$ due to the delay imposed in following these paths via self-loops even though the associated shortest paths might have common states. This and (74) imply that Condition $(i)$ in Lemma 1 holds. This completes the proof of Claim 1. ∎
Claim 2.
The discrete processes $\{\widehat{w}_{k}\}$ and $\{(w_{k},z_{k})\}$ satisfy Condition $(ii)$ in Lemma 1 with $\lambda_{(n)}=1-CHK_{(n)}\times\beta_{(n)}$, for some constant $C>0$, $n$ is sufficiently large.
Proof.
The transition probabilities $\mathbb{P}^{w}(\cdot\mid\{w_{l},z_{l}\}_{l\leq k})$ and ^$\widehat{\mathbb{P}}(\cdot\mid\widehat{w}_{k})$ are linear functions of the softmax response, resp., to the Q-function estimates $Q_{t_{k+1}}(\cdot)$ and $Q_{(n)}(\cdot)$. Particularly, the transition of the current action profile depends on the agents’ softmax response if they are updating their actions and agents update their actions with certain probability either in a coordinated way as in coordinated scheme or in an independent way as in independent scheme. Furthermore, the softmax response (6) is a $1/\tau$-Lipschitz continuous function of the Q-function estimates with respect to $\|\cdot\|_{2}$ (Gao and Pavel, 2018, Proposition 4). Since norms are comparable in finite-dimensional spaces, there exists $C>0$ such that for any $k$, action profile $a$, and trajectory $\{z_{l}\}_{l=k_{(n)}}^{k}$, we have
^$\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\|\widehat{\mathbb{P}}(\cdot\mid w_{k})-\mathbb{P}^{w}(\cdot\mid\{w_{l},z_{l}\}_{l\leq k})\right\|_{\mathrm{TV}}\!\!$ $\leq C\|Q_{t_{k+1}}-Q_{(n)}\|_{\max}\leq CHK_{(n)}\beta_{(n)},$ (75)
for all $k\in[k_{(n)},k_{(n+1)})$, where $\beta_{(n)}:=\max_{a}\{\beta_{c_{(n)}(s,a)}\}$, and the last inequality follows from the boundedness of the Q-function estimates and the monotonicity of the step size. Set $\lambda_{(n)}=1-CHK_{(n)}\times\beta_{(n)}$. Then, $\lambda_{(n)}$ takes a value in $(0,1]$ for sufficiently large $n$ since $K_{(n)}\times\beta_{(n)}$ is decaying to zero as $n\rightarrow\infty$ by Lemma 4 whereas $C$, and $H$ are time-invariant. Hence, Condition $(ii)$ in Lemma 1 also holds for sufficiently large $n$. This completes the proof of Claim 2. ∎
By Claims 1 and 2, Lemma 1 yields that
$\|\mu_{k}^{w}-\widehat{\mu}_{k}\|_{1}\leq 2(1-\varepsilon)^{\frac{(k-k_{(n)})}{\kappa}-1}+2\left(1-\lambda_{(n)}^{\kappa}\right)\frac{1+\varepsilon}{\varepsilon}$ (76)
for all $k\geq k_{(n)}$. Recall that $\{\widehat{w}_{k}\}$ forms an irreducible and aperiodic Markov chain. Let $\widehat{\mu}_{(n)}\in\Delta(W)$ denote its unique stationary distribution, i.e., $\widehat{\mu}_{k}\rightarrow\widehat{\mu}_{(n)}$ as $k\rightarrow\infty$. Since $\widehat{\mu}_{(n)}$ is the stationary distribution of the Markov chain $\{\widehat{w}_{k}\}_{k\geq k_{(n)}}$, if we had $\widehat{\mu}_{k_{(n)}-1}=\widehat{\mu}_{(n)}$, i.e., $\widehat{a}_{k_{(n)}}\sim\widehat{\mu}_{(n)}$, then we would have $\widehat{\mu}_{k}=\widehat{\mu}_{(n)}$, i.e., $\widehat{w}_{k}\sim\widehat{\mu}_{(n)}$, for all $k\geq k_{(n)}$. Therefore, (76) yields that
$\|\mu_{k}^{w}-\widehat{\mu}_{(n)}\|_{1}\leq 2(1-\varepsilon)^{\frac{(k-k_{(n)})}{\kappa}-1}+2\left(1-\lambda_{(n)}^{\kappa}\right)\frac{1+\varepsilon}{\varepsilon}$ (77)
for all $k\geq k_{(n)}$ which completes the proof of Lemma 2.
Appendix C Proof of Lemma 3
The proof is by construction. First of all, by (33), we have
$\alpha_{k_{(n+1)}-1}\leq\overline{\alpha}_{(n)}=\alpha_{k_{(n)}}\prod_{l=k_{(n)}+1}^{k_{(n+1)}-1}(1-\alpha_{l})+\ldots+\alpha_{k_{(n+1)}-1}\leq K_{(n)}\alpha_{k_{(n)}}$ (78)
since $\alpha_{k}$’s are monotonically non-increasing and take values in $[0,1]$. As $\alpha_{k}$ decays to zero, the bounds in (78) would imply that $\overline{\alpha}_{(n)}$ also decays to zero if we had $\lim_{n\rightarrow\infty}K_{(n)}\alpha_{k_{(n)}}=0$ which is true if $K_{(n)}$ grows at a sub-linear rate, and $k_{(n)}\geq n$. Some algebra would also show that $\overline{\alpha}_{(n)}=1-\prod_{k=k_{(n)}}^{k_{(n+1)}-1}(1-\alpha_{k})$, and therefore, $\overline{\alpha}_{(n)}\in[0,1]$. Next we show that such a realization of $K_{(n)}$ is possible such that $\overline{\alpha}_{(n)}$ satisfies standard conditions.
Define the following:
$\overline{A}_{N}\coloneqq\sum_{k=1}^{N}\alpha_{k},\quad\quad\underline{A}_{N}\coloneqq\sum_{k=N}^{\infty}\alpha_{k}^{2}.$ (79)
By definition and the assumptions stated in the Lemma 3, $\lim_{N\rightarrow\infty}\overline{A}_{N}=\infty$, and $\lim_{N\rightarrow\infty}\underline{A}_{N}=0$. Then, for any $m\in\mathbb{N}$,
$\overline{\nu}(m)\coloneqq\min\{N:\overline{A}_{N}\geq 2^{m},N\in\mathbb{N}\},\quad\underline{\nu}(m)\coloneqq\min\{N:\underline{A}_{N}\leq 2^{-m},N\in\mathbb{N}\}$ (80)
are well defined functions. $\overline{\nu}(m)$ and $\underline{\nu}(m)$ are monotonically non-decreasing by definition. Also, define the following:
$\overline{K}_{(n)}=\max(1,\max\{m:\overline{A}_{n}\geq 2^{m},m\in\mathbb{N}\}),$ (81)
$\underline{K}_{(n)}=\max(1,\max\{m:\underline{A}_{n}\leq 2^{-m},m\in\mathbb{N}\}).$ (82)
Note that, $\overline{K}_{(n)}=m$ if and only if $\overline{\nu}(m)\leq n<\overline{\nu}(m+1)$, and $\underline{K}_{(n)}=m$ if and only if $\underline{\nu}(m)\leq n<\underline{\nu}(m+1)$. Based on these definitons, we let $K_{(n)}=\min(\overline{K}_{(n)},\underline{K}_{(n)})$. Notice that, by this definition $K_{(n)}$ grows at a sub-linear rate.
Due to the epoch length having a minimum length of $1$, i.e. $K_{(n)}\geq 1$, and $k_{(1)}=1$, we have $k_{(n)}\geq n$. Morever, since $\alpha_{k}\in[0,1)$, we have $\overline{\nu}(m)\geq m$ for all $m$.
By the left inequality in (78), we have
$\sum_{n=1}^{\infty}\overline{\alpha}_{(n)}\geq\sum_{n=1}^{\infty}\alpha_{k_{(n+1)}-1}$ $\geq\sum_{n=1}^{\infty}\sum_{k=k_{(n+1)}-1}^{k_{(n+2)}-1}\frac{\alpha_{k}}{K_{(n+1)}}$ (83)
$\geq\sum_{k=k_{(2)}-1}^{\infty}\frac{\alpha_{k}}{K_{(k+1)}}$ (84)
$\geq\sum_{k=k_{(2)}-1}^{\infty}\frac{\alpha_{k}}{\overline{K}_{(k+1)}}$ (85)
$\geq\sum_{m=k_{(2)}}^{\infty}\sum_{k=\overline{\nu}(m)}^{\overline{\nu}(m+1)-1}\frac{\alpha_{k}}{\overline{K}_{(k+1)}}$ (86)
$\geq\sum_{m=k_{(2)}}^{\infty}\frac{1}{m+1}\left(\sum_{k=1}^{\overline{\nu}(m+1)}\alpha_{k}-\sum_{k=1}^{\overline{\nu}(m)-1}\alpha_{k}-1\right)$ (87)
$\geq\sum_{m=k_{(2)}}^{\infty}\frac{1}{m+1}(2^{m+1}-2^{m}-1)$ (88)
$\geq\sum_{m=k_{(2)}}^{\infty}\frac{2^{m}-1}{m+1}=\infty.$ (89)
By the right inequality in (78), we have
$\sum_{n=\underline{\nu}(1)}^{\infty}\overline{\alpha}_{(n)}^{2}\leq\sum_{n=\underline{\nu}(1)}^{\infty}K_{(n)}^{2}\alpha_{k_{(n)}}^{2}\leq\sum_{n=\underline{\nu}(1)}^{\infty}\underline{K}_{(n)}^{2}\alpha_{n}^{2}$ $\leq\sum_{m=1}^{\infty}\sum_{n=\underline{\nu}(m)}^{\underline{\nu}(m+1)-1}\underline{K}_{(n)}^{2}\alpha_{n}^{2}$ (90)
$\leq\sum_{m=1}^{\infty}(m)^{2}\sum_{n=\underline{\nu}(m)}^{\underline{\nu}(m+1)-1}\alpha_{n}^{2}$ (91)
$\leq\sum_{m=1}^{\infty}(m)^{2}\sum_{n=\underline{\nu}(m)}^{\infty}\alpha_{n}^{2}$ (92)
$\leq\sum_{m=1}^{\infty}\frac{(m)^{2}}{2^{m}}\leq\infty.$ (93)
For the claim that $K_{(n)}/k_{(n)}$ decaying to zero, we can simply use the definition of $K_{(n)}$. Since $\alpha_{k}<1$ for all $k$, we can say that $\overline{A}_{2^{m}}<2^{m}$. Then, $\overline{K}_{(2^{m})}<m$. Since $\{\overline{K}_{(n)}\}_{n\geq 1}$ is a monotonically non-decreasing sequence, $\overline{K}_{(n)}\leq\lceil\log_{2}(n)\rceil$. Also, we know that $k_{(n)}\geq n$. Then, the following is true
$0\displaystyle 0\leq\lim_{n\rightarrow\infty}\frac{K_{(n)}}{k_{(n)}}\leq\lim_{n\rightarrow\infty}\frac{\overline{K}_{(n)}}{n}\leq\lim_{n\rightarrow\infty}\frac{\lceil\log_{2}(n)\rceil}{n}=$ 0 . (94)
This completes the proof.
Appendix D Proof of Proposition 3
Let $a_{t}\in A$ and $\underline{a}_{t}\in A$ denote the action profiles, resp., in the classical and independent log-linear dynamics. Correspondingly, $\mu_{t}\in\Delta(A)$ and $\underline{\mu}_{t}\in\Delta(A)$ denote, resp., the distribution over $a_{t}$ and $\underline{a}_{t}$. Both $\{a_{t}\}_{t\geq 0}$ and $\{\underline{a}_{t}\}_{t\geq 0}$ form irreducible Markov chains with self-loops at every state.
Let $P$ and $\underline{P}$ denote the transition matrices in the Markov chains, resp., associated with the classical and independent log-linear learning dynamics. The ways agents update their actions in both dynamics yield that the transition matrices $P$ and $\underline{P}$ satisfy $\underline{P}=p_{0}I+p_{1}P+(1-p_{0}-p_{1})E$, where $p_{0}:=(1-\delta)^{n}\in(0,1)$ and $p_{1}:=n\delta(1-\delta)^{n-1}\in(0,1)$ correspond, resp., to the probability that none and one of the agents update their actions, and $E$ is the transition matrix corresponding to the case at least two (randomly picked) agents updates their actions. The transition matrix corresponding to the case that at least one (randomly picked) agent updates their actions is given by
$\overline{P}=\frac{p_{1}}{1-p_{0}}P+\frac{1-p_{0}-p_{1}}{1-p_{0}}E.$ (95)
Notably, the transition matrix $\overline{P}$ induces an irreducible and aperiodic Markov chain whose unique stationary distribution is also $\underline{\mu}$, i.e., $\overline{P}\underline{\mu}=\underline{\mu}$ as $\underline{P}\underline{\mu}=\underline{\mu}$. Since $\overline{P}$ and $\underline{P}$ induce the same unique stationary distribution, we focus on the processes $\{\overline{a}_{t}\}$ evolving according to $\overline{P}$ rather than $\underline{P}$, and couple the evolution of $\{\overline{a}_{t}\}$ and $\{a_{t}\}$ based on Lemma 1.
Given some $a^{\prime}\in A$, consider any pair $(a,\underline{a})\in A\times A$ where $a\neq\underline{a}$, e.g., $a^{j}\neq\underline{a}^{j}$ for some $j\in[N]$. Recall that the update of actions to any action at most for one agent is feasible for both coordinated and independent log-linear learning dynamics. Therefore, there exists $n$-length paths $a=h^{0},\ldots,h^{n}=a^{\prime}$ and $\underline{a}=\underline{h}^{0},\ldots,\underline{h}^{n}=a^{\prime}$; where consecutive action profiles $(h^{k},h^{k+1})$ and $(\underline{h}^{k},\underline{h}^{k+1})$ differ at most for one agent’s action, and the action profiles $\{h^{k}\}_{k=0}^{n-1}$ and $\{\underline{h}^{k}\}_{k=0}^{n-1}$ have, resp., the same actions $a^{j}$ and $\underline{a}^{j}$ for agent $j$. This implies that these paths do not visit the same action profile concurrently for $k=0,\ldots,n-1$. Furthermore, the processes $\{a_{t}\}$ and $\{\underline{a}_{t}\}$ can follow these paths with probabilities bounded from below by
$\Pr\left(\bigwedge_{k=1}^{n}a_{t+k}=h^{k}\mid a_{t}=h^{0}\right)\geq\left(\frac{\epsilon}{n}\right)^{n},$ (96a)
$\Pr\left(\bigwedge_{k=1}^{n}\underline{a}_{t+k}=\underline{h}^{k}\mid\underline{a}_{t}=\underline{h}^{0}\right)\geq\left(\frac{\epsilon\delta(1-\delta)^{n-1}}{1-(1-\delta)^{n}}\right)^{n}.$ (96b)
Therefore, Condition $(i)$ in Lemma 1 holds.
On the other hand, (95) yields that
$\left\|P(\cdot\mid h)-\overline{P}(\cdot\mid h)\right\|_{\mathrm{TV}}$ $=\frac{1-p_{0}-p_{1}}{1-p_{0}}\cdot\left\|P(\cdot\mid h)-E(\cdot\mid h)\right\|_{\mathrm{TV}}$ (97)
for any $h\in A$. Note that $\left\|P(\cdot\mid h)-E(\cdot\mid h)\right\|_{\mathrm{TV}}\leq 1$. Therefore, Condition $(ii)$ in Lemma 1 holds for $\lambda=1-(1-p_{0}-p_{1})/(1-p_{0})=p_{1}/(1-p_{0})\leq 1$. Since Conditions $(i)$ and $(ii)$ hold, we can invoke Lemma 1 as let $t\rightarrow\infty$ to conclude (25).